### v1

# ObjectFinder: Open-Vocabulary Assistive System for Interactive Object Search by Blind People

Abstract.
Assistive technology can be leveraged by blind people when searching for objects in their daily lives. We created ObjectFinder, an open-vocabulary interactive object-search prototype, which combines object detection with scene description and navigation. It enables blind persons to detect and navigate to objects of their choice. Our approach used co-design for the development of the prototype. We further conducted need-finding interviews to better understand challenges in object search, followed by a study with the ObjectFinder prototype in a laboratory setting simulating a living room and an office, with eight blind users. Additionally, we compared the prototype with BeMyEyes and Lookout for object search. We found that most participants felt more independent with ObjectFinder and preferred it over the baselines when deployed on more efficient hardware, as it enhances mental mapping and allows for active target definition. Moreover, we identified factors for future directions for the development of object-search systems.
People with visual impairments, Assistive technology,
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Accessibility systems and toolsccs: Computing methodologies Computer vision tasks
1. Introduction
We all regularly visit unfamiliar buildings and shared spaces, and need to deal with uncertainty about our environment. This also includes blind people  (Müller et al., 2022; Zhang et al., 2021), for whom rapidly changing environments can be challenging to navigate, e.g., finding the way around a room, detecting unexpected obstacles (Alkhanifer and Ludi, 2015; Jeamwatthanachai et al., 2019) and searching for objects without that may have been moved by others. When first entering a new or changed environment, it is important to get an overview of the arrangement of furniture, sanitary facilities, and room installations in order to build up a mental map of a room (Façanha et al., 2020). Here, commonly leveraged haptic exploration of the environment can have safety implications due to the need for physical contact with potentially hazardous objects or surfaces (Zheng et al., 2024; Gonzalez Penuela et al., 2024), such as knifes and hot drinks. Assistive technology has attempted to fill this gap, with some applications (Eyes, 2024; Aira Company, 2024) connecting blind individuals with remote volunteers, enforcing reliance on non-disabled persons for assistance. Other approaches leverage AI  (Eyes, 2024; Achiam et al., 2023; Corporation, 2024; Lee et al., 2022a; Liu et al., 2023) to facilitate exploration of unfamiliar environments.
Such technologies can be divided into two categories: description-based and detection-based systems. Description-based systems provide either a vivid and detailed descriptions of a photo (Eyes, 2024) or brief captions (Corporation, 2024; Lee et al., 2022a; Liu et al., 2023). However, these systems are not able to localize a specific object in an unfamiliar environment while ensuring that the target object is in the image area, as illustrated on the left side of Figure 1 (Challenge 1, C1). As far as detection-based systems are concerned (Google, 2024; V7, 2024; Zheng et al., 2024), current technological approaches only allow the search for specific objects (Constantinescu et al., 2020a; Yi et al., 2013; Schauerte et al., 2012; Ahmetovic et al., 2020; Kacorri et al., 2017; Google, 2024; V7, 2024). Therefore, when using current detection-based systems in an unfamiliar environment, blind users may not know what there is in a room and miss unexpected items that could be of interest, such as a plate of cookies in an unknown office, as shown on the right side of Figure 1 (Challenge 2, C2).
To address these challenges, we designed ObjectFinder, an interactive prototype for open-vocabulary object search by combining both description-based and detection-based methods. The integration of Open-Vocabulary Object Detectors (OVOD) (Cheng et al., 2024; Li et al., 2024) and Multimodal Large Language Models (MLLM) (Achiam et al., 2023; Liu et al., 2024) in ObjectFinder makes it possible to develop an interactive process ranging from object detection to description: input any target to get detailed information about the surrounding region of interest. This, in turn, identifies potentially interesting and unexpected targets, which can then be searched for in subsequent loops.
Through our work, we seek to answer the following two research questions:
  • RQ1: How can an open-vocabulary prototype integrate detection-based and description-based technologies to support object search for blind people, and what challenges are associated with its design and development?
  • RQ2: What are users’ experience and preferences regarding the open-vocabulary interactive object search system, specifically comparing the description-based and the detection-based, with the combined (ObjectFinder) systems?
We structured our work into three phases. In the first phase, we co-designed with a blind person throughout the conception and development of our system and study design. Thus, we first identified a use case that includes the following functions: target object detection (F1), localization (F2), route planning (F3), scene description (F4), and open questions (F5). Since existing description-based and detection-based systems are not tailored for object-search tasks, we developed a competitive baseline, ObjectFinder_Base, with limited detectable objects. Then, we conducted four half-hour user-in-the-loop sessions with our blind co-designer, inviting him to explore features in different scenarios (a living room and an office), leading to the refinement of ObjectFinder. In the second phase, we conducted a semi-structured interview with eight blind participants to understand their daily targets of object search with their currently used methods, and challenges they encounter in unfamiliar or changing environments. In the third phase, we conducted an exploratory evaluation of ObjectFinder inviting the same eight participants to engage with the system, and take part in a semi-structured interview afterwards. Through thematic analysis (Fereday and Muir-Cochrane, 2006), we show that the use of ObjectFinder facilitated interactive object search because it combines crucial information about localization (distance and direction) through the detection part with relationships among objects through the description part, as well with the help in navigating to the searched objects. However, we have also found that the current system is not efficient enough, is perceived too cumbersome and is not available on the cell phone, and thus cannot be integrated into daily live.
Our work makes the following key contributions: (1) We make a system contribution and introduce an open-vocabulary interactive object-search prototype that supports location of objects and provision of relevant information. (2) We provide an overview of target objects and search challenges in the daily lives of participants under consideration of the potential of assistive technology. (3) We critically reflect on the experience and effectiveness of our prototype, and outline opportunities and tensions to be considered in the future development of open-vocabulary interactive object-search systems for real-world application.
2. Related Work
In this section, we provide an overview of existing description-based and detection-based systems for scene understanding and exploration. We also discuss the integration of multimodal large language models and interaction features in assistive technologies, which form the background of our study.
2.1. Description-based and Detection-based Systems for Blind People
With the advancement of Artificial Intelligence (AI), researchers are developing supplementary AI-based systems to assist blind people in exploring scenes. These systems can be classified into description-based and detection-based. The description-based systems capture the scene and describe it only once. Seeing AI (Corporation, 2024), ImageExplorer (Lee et al., 2022a) and OpenSU (Liu et al., 2023) generate brief image captions for the scenes captured by a mobile phone and enable tactile exploration of the salient objects on the touchscreen. TapTapSee (TapTapSee, Inc., nd) is an application that generates a concise phrase about the salient object in almost real-time. Be My AI, a feature of Be My Eyes (Eyes, 2024) empowered by GPT-4, delivers vivid descriptions of the scenario and allows users to ask questions. Some other works focus on the specific features of the salient object, e.g., material (Zheng et al., 2024), transparency (Zhang et al., 2021), and various hazards (Yang et al., 2018a, b). However, people with blindness should make the object within the region of interest captured by mobile devices while using description-based applications (Yang et al., 2024; Gurari et al., 2018). In addition, we found that people with blindness can barely align the photos they capture with real spatial dimensions using mobile phones during our user study which is in line with the findings of (Kacorri et al., 2017). Therefore, we implemented an object detector to identify the region of interest for the description module, ensuring a precise understanding of the area where the target is located.
Detection-based systems are designed to provide real-time outputs of identified objects or features of interest. Lookout (Google, 2024) and AIPoly (V7, 2024) exemplify this capability by identifying the nearest object within the phone’s field of view. Various studies have developed wearable systems (Liu, 2023; Islam et al., 2023; Sugashini and Balakrishnan, 2024) with similar functionalities, offering real-time object information through multiple interaction modes. Research has explored the detection of personal objects using methods such as SIFT (Constantinescu et al., 2020a; Yi et al., 2013; Schauerte et al., 2012) and advanced deep learning networks (Ahmetovic et al., 2020; Kacorri et al., 2017). Navigation assistive systems utilize detection-based methods for obstacle avoidance (Bala et al., 2023; Liu et al., 2021a; Ou et al., 2022), risk assessment (Wang et al., 2024), object finding (Li et al., 2023; Duh et al., 2020; Hu et al., 2022) and passable path planning (Zou et al., 2023; Surougi and McCann, 2023; Hong et al., 2024; Jain et al., 2023). These systems autonomously select information, which may limit user agency in actively defining targets. Constantinescu et al. (Constantinescu et al., 2020b) propose a system that allows users to choose from a limited set of objects to receive audible feedback in their vicinity. In detection-based systems, the information available to users is constrained by system limitations, preventing awareness of all existing objects in the environment. This study integrates the strengths of description-based and detection-based systems to identify regions of interest and inspire users to unexpected targets.
Figure 2. Use case diagram : A blind person uses a system to find a socket in an unfamiliar office. The step with dashes is not preferred.
[…]
There are eight steps in the diagram: 1) The user tells the system what he/she is looking for (e.g., a socket). 2)The system gives feedback that it understands what the user wants. 3)The system indicates it is ready for searching (through a sound). 4)The user looks around using the system. (Function 1) 5a) The system only announces sockets. 5b) The system continuously informs the user of all objects in the frame (e.g., chairs, table). 6) Socket detected. The system provides feedback with the distance, direction, and relevant points of interest (POls) in the surroundings. (Function 2, Function 4). 7a) The user wants to continue searching. The system keeps searching without intervention. 7b) The user is happy with the found object. He/she asks for further details about it, such as directions on how to reach the object and the height of the socket on the wall. (Function 5). 8) For large rooms, the system should navigate the user (e.g., ”Walk around the table on the left side”). (Function 3)
Figure 3. Hardware design and components of the wearable system ObjectFinder. It incorporates a stereo camera to capture visual information about the user’s surroundings, a pair of buttons and a microphone to collect the user’s commands. Simultaneously, it executes algorithms through a lightweight processor. To provide a comprehensive and immersive experience, the system delivers spatial-aware informational feedback directly to the user via bone-conduction headphones.
[…]
This image displays the components and connectivity of a wearable system. The system consists of smart glasses equipped with a stereo camera, a microphone on the collar, and two buttons on the waist bag for user inputs. The smart glasses also include bone conduction earphones for output. The image further details the internal components of the glasses which are magnified to show the arrangement from right to left: the right earphone, right infrared camera, color camera, left infrared camera, image ASIC, and left earphone. The NVIDIA Jetson Nano, housed in the waist bag, processes the data collected by these components.
2.2. Multimodal Large Language Models for Assistive Technology for Blind People
Be My Eyes (Eyes, 2024) is the pioneering application to collaborate with OpenAI, integrating GPT-4 (Achiam et al., 2023) to enable blind individuals to comprehend the real world and interpret images. VIALM (Zhao et al., 2024) evaluates the existing Large Language Models (LLMs) for navigation instructions, identifying GPT-4 as the state-of-the-art based on both automatic and human evaluation metrics. VIAssist (Yang et al., 2024) identifies six challenges in capturing desirable images for daily requests, such as incomplete target object placement, and offers reshoot suggestions for low-quality images. VQAsk (De Marsico et al., 2024) is an Android application that enables interaction with one’s photographs or the surrounding environment through speech and haptic feedback. VisionGPT (Wang et al., 2024) leverages LLM to identify existing objects in different scenarios and assess the risk. nanoLLaVA (Face, nd) is the existing smallest LLM developed for the deployment on edge devices, e.g. assistive technology for blind people. Waisberg et al. (Waisberg et al., 2024) proposed embedding LLMs in Ray-Ban glasses from Meta to capture the egocentric view of blind people. To enable blind people to access visual arts, Zhang et al. (Zhang, 2024) designed an interface that enables blind individuals to access visual information from images using LLMs and VLMs. GenAssist (Huh et al., 2023) makes image generation accessible by utilizing a sequence of prompts for LLMs. In this work, we develop our ObjectFinder by using GPT-4 to generate the description and answer questions.
In order to generate proper information from LLM, OpenAI offers an extensive tutorial (OpenAI, 2023) on strategies for crafting effective prompts to achieve optimal results. Numerous studies have explored various aspects of prompt engineering. In this context, we present specific prompt engineering projects designed for the blind community. VIALM (Zhao et al., 2024) identifies the importance of prompting LLMs to generate detailed information regarding obstacles and tactile features in navigation instructions. Gadiraju et al. (Gadiraju et al., 2023) investigate how LLMs can perpetuate subtle yet harmful biases towards communities with special needs, identify specific categories of these biases, and propose training models on diverse resources that positively represent individuals with special needs to mitigate such harms. Hao et al. (Hao et al., 2024) customize prompts for blind people for various purposes and eliminate any potential offense. In this work, we crafted prompts using these strategies to enhance interactive interfaces for blind people, improving digital accessibility and inclusivity.
2.3. Interaction Features
Interaction features are crucial for enhancing the usability and effectiveness of assistive technology for blind persons. Here we discuss the in- and output modalities separately. Speech command input provides a more intuitive and human-like interaction without physical touch, exemplified by voice-controlled smart home devices that offer safety and convenience for blind people (Lee et al., 2022b). Voice assistants such as Apple’s Siri and Amazon’s Alexa are widely utilized by the blind community and are deemed beneficial (Oumard et al., 2022b). Moreover, advanced navigation systems employ voice commands to define targets and utilize GPS for localization (Hersh, 2022; Ramadhan, 2018). He et al. (He et al., 2023) enhance the speech command with bone-conducted vibration on head-mounted wearables. However, the efficacy of speech understanding systems is contingent upon high accuracy, which can be significantly impacted by environmental noise and the systems’ limited ability to comprehend diverse vocabularies (Oumard et al., 2022a). To address these issues, some solutions incorporate simple button-based inputs as alternatives to voice commands (Awad et al., 2018; Tahoun et al., 2020; Hersh, 2022). Kamikubo et al. (Kamikubo et al., 2024) found that in their study, most users referred to active methods of using buttons, and few users prefer to use voice commands to ask specific questions. PathFinder (Kuribayashi et al., 2023) is an automated navigation system with four buttons on the handle interface for functionality selection. In our work, we utilize brief voice commands to enable flexible definition of target objects, while employing buttons to streamline the selection of functions.
In the realm of assistive technologies, methods for conveying information serve distinct purposes. Haptic cues, noted for their intuitive reliability, prove effective across indoor and outdoor settings (Bouteraa, 2021). Vibration is widely employed for alerts in wearable devices (Mahmud et al., 2014; Babu D et al., 2021; Bouteraa, 2021), while tactile cues (Pielot et al., 2012; Liu et al., 2021b) and kinetic force (Amemiya and Sugiyama, 2008, 2009; Ando et al., 2012) are utilized for indicating navigation directions with handheld devices. Yang et al. (Yang et al., 2021) explore light cues on hat brims to guide individuals who have light perception. Auditory signals, such as beeps of varying frequencies, play crucial roles in warning systems for users and pedestrians (Guerrón et al., 2020; Kayukawa et al., 2019). Earcons, which provide contextual information about navigated interfaces, enhance user understanding (Perkins School for the Blind, 2023; OrCam Technologies, 2022). Speech remains the predominant output feature on mobile phones (Leporini et al., 2012; Rodrigues et al., 2015) and wearable systems (Bouteraa, 2021; Chen et al., 2021; Boldu et al., 2020; Liu et al., 2021a; Li et al., 2023), offering comprehensive information delivery. Notably, bone-conduction devices are preferred for their comfort and wearability, enabling the hearing of system instructions alongside environmental sounds (Asakura, 2021; Liu et al., 2021a). Our approach integrates earcons and speech output via bone-conduction headphones within a pair of smart glasses to enhance information accessibility.
In the remainder of this paper, we outline our three-stage research process for developing our object search prototype: (1) a user-in-the-loop approach for designing and refining the prototype, (2) conducting interviews with additional participants to understand their target objects and challenges related to object search, and (3) performing exploratory user studies in both public and private settings to compare the prototype with existing description-based and detection-based systems.
3. Phase I: User-in-the-loop Design
In this phase, we engaged in one-on-one co-design with a blind person (P0, see Table 1) to define functions of the first prototype of the ObjectFinder system, and to develop the design of a study with potential users of the system, rooting initial design decision in the lived experience of a blind person. In this phase, we focus on the challenge of combining description-based and detection-based systems as a foundation to address RQ1. Here, we provide an overview of our design approach, detail our prototype development process, and present the resulting system.
Figure 4. Left: The initial flowchart based on use case definition. The text inside the quotation marks (e.g., “Find”) represents voice input keywords to trigger target identification, target confirmation, and function selection. Right: The system architecture includes the interaction steps and multiple functions corresponding to the flowchart. Given a task (e.g., “finding a coffee table in an unknown space”), a target with location and orientation information can be found and output. The visualization is derived from a real example in our user study.
[…]
This figure illustrates the initial flowchart based on the use case definition on the left side and the system architecture on the right side. The flowchart has the following elements: 1. Vocal input: ’Find ¡target¿’ (block circled 1) a. Flows to 2. 2. Confirm ¡target¿? a. If yes, flows to 3. b. If no, flows back to 1. 3. Scan the environment, while continuously judging if the target is found. (block circled 2) a. If yes, flows to 4. 4. Localization information: ¡target¿ ¡distance¿ ¡direction¿ (block circled 3) a. Flows to 5. 5. Voice command to choose route planning, scene description, or open questions. (block circled 4) a. If the voice command ’continue searching’ is given, flows back to 1. On the right side of the figure, the network architecture, including the inputs and outputs of YOLO World and MobileSAM, is depicted. These components are illustrated to demonstrate their functionality in efficient object detection and enhancing the precision of localization information.
3.1. Definition of the Use Case
In a first step, we started with the use case definition, following the principle in (Cockburn, 2000). For this purpose, we organized a workshop that included the blind co-designer (P0), one developer, and two accessibility and usability experts (one of them is a blind person). The use case search for an object, here a socket, in an unfamiliar environment was identified in the workshop. Then, the blind co-designer had the task imagining how he would use a system to look for a socket in an unknown office, while the other participants asked questions that helped the co-designer to define the requirements and the different steps of the search. We formulated the use case in a diagram in Figure 2 for a clear illustration. The co-designer’s expectation suggested decomposition of the use case into $5$ functions: Object Detection (F1), Localization (F2), Route Planning (F3), Scene Description (F4), and Open Questions (F5). Leveraging these categories, we first restructured the interaction features and functions into an initial flowchart, as shown in Figure 4. In the following, we explain assembly of hardware components of ObjectFinder, and we outline implementation of basic functionality into the first prototype.
3.2. Iteration 1: Initial Prototype Design
In this section, we describe the implementation of the initial prototype based on this use case, detailing the hardware and interaction features and the functionality implementation.
3.2.1. Hardware and Interaction Features
Figure 3 presents the system diagram, which comprises a pair of KRVision (KR Vision, [n. d.]) smart glasses coupled with a waist bag. The smart glasses are outfitted with a RealSense R200 RGB-Depth camera, facilitating real-time RGB and depth frame acquisition in an egocentric perspective. Additionally, a bone conduction headset is incorporated, enabling auditory output while maintaining perception of environmental sounds. In the waist bag, a Jetson Nano, a compact and powerful processor, is utilized for efficient data processing, accompanied by a power bank for energy supplementation. The waist bag features two buttons which can be programmed and used for volume adjustment initially. A microphone is attached to the collar for audio input. As shown in Figure 4, participants simply say commands with keywords to trigger target identification, target confirmation, and function selection.
Figure 5. Rectified flowchart. Left: The optimized version of the flowchart for the open-vocabulary interactive ObjectFinder. Right: Steps added to the ObjectFinder baseline. The object detection model in the baseline can only identify a list of objects, the added steps assist in locating the other objects through egocentric scene descriptions.
[…]
On the left side of the figure is the rectified flowchart, and on the right side are the different steps between ObjectFinder and ObjectFinder_Base. The rectified flowchart includes the following steps: 1. Vocal input: ’Find ¡target¿’ (block circled 1) a. Flows to 2. 2. Confirm ¡target¿? a. If yes, flows to 3. b. If no, flows back to 1. 3. Scan the environment while continuously judging if the target is found. (block circled 2) a. If yes, flows to 4. b. If time out, flows to 5. 4. Localization information: ¡target¿ ¡distance¿ ¡direction¿ (block circled 3) a. Flows to 6. 5. Press 1 and 2 to trigger continued searching and scene description. (block circled 5) a. If 1 is pressed, flows back to 1. 6. Press 1, 2, 3, and 4 to trigger route planning, scene description, continued searching, and open questions. (block circled 4) a. If 1 is pressed, flows back to 1. b. Otherwise, conduct the function and flow to 7. 7. Press 1, 2, and 3 to trigger continued searching, repeat, and either scene description or open questions. (block circled 6) a. If 1 is pressed, flows back to 1. b. Otherwise, remain in this block. On the right, the different steps between blocks circled 1 and 2 are illustrated: 1. Confirm ¡target¿? a. If yes, flows to 2. 2. Judge if the target is in COCO2017. a. If yes, flows to circled 2. b. If no, flows to 3. 3. Press 1 to continue searching, press 2 for scene description.
3.2.2. Functionality Implementation
To achieve the aforementioned open-vocabulary interaction, all five functions are primarily realized by a combination of models: YOLO-World (Cheng et al., 2024) and MobileSAM (Zhang et al., 2023) as Open-Vocabulary Object Detection (OVOD), GPT-4 (Achiam et al., 2023) as Multimodal Large Language Model (MLLM). As the initial flowchart and architecture of the system are shown in Figure 4, we detail the implementation of the individual function in the following.
Object Detection (F1).
The recently proposed YOLO-World (Cheng et al., 2024), employing a prompt-then-detect strategy, allows users to define their customized vocabularies for the object detector. In our prototype, users specify the target object via a voice command “Find <target>”. The prototype then verifies whether the object exists within the predefined vocabulary list. If absent, the new target word is embedded and saved in the model’s vocabulary for future reference, a process requiring approximately $20$ seconds for initialization. To enhance efficiency, we have pre-populated the vocabulary with a list of common everyday objects. We utilize the large version of YOLO-World, which performs inference in $0.5$ seconds on our system. The key frame where the object is detected is retained for subsequent analysis.
Localization (F2).
Our study utilized an Object-Distance-Direction format for localization as proposed by Constantinescu et al. (Constantinescu et al., 2022). Directions are presented in a clockwise orientation and distances in meters. Initial distances were estimated using YOLO-World’s average bounding box distances due to the pending release of its segmentation version. These bounding boxes, which need to overlap ground truth by at least $50\%$, might inaccurately include significant background areas in calculations (Liu et al., 2023). To enhance accuracy, MobileSAM (Zhang et al., 2023), the compact version of SAM (Kirillov et al., 2023), was later added to generate segmentation masks using the bounding boxes as prompts post YOLO-World (Cheng et al., 2024).
Multimodal Large Language Model Functions (F3-F5).
According to VIALM (Zhao et al., 2024), GPT-4 is the state-of-the-art for guiding blind people, excelling in both human (correctness, actionability, fluency) and automated evaluations. Thus, we chose GPT-4 for implementing MLLM functions (F3-F5). Effective prompt engineering is crucial for enhancing large language models’ utility and accessibility for blind users. Our system prompt is concrete, incorporating role, tone, and response length, guided by OpenAI’s Prompt Engineering tutorial (OpenAI, 2023). Response length, aligned with social media alt-text limits (100-500 characters) (for the Blind, 2024), is set to the maximum to cater to user preferences for vivid responses. For navigation instructions, Wang et al. (Wang et al., 2023) suggest including position, count, size, color, material, and shape. We developed two user prompts: one for route planning (F3) and another for scene description (F4), with voice input for open questions (F5). The concrete prompts are provided in the supplementary material.
3.3. Iteration 2: System Optimization and Refinement
We conducted four rounds of preliminary half-hour user exploration, inviting P0 to explore the prototype freely in a demo room that integrates two scenarios: a private setting (i.e., a living room) and a public setting (i.e., an office). The layouts of these scenarios are shown in Figure 7(b). P0 could search for anything and trigger any functions he wanted. The wearable prototype is designed to free the hands of blind people to hold a white cane. However, it was difficult to determine which information was acquired from our prototype in the small scenarios. Therefore, we decided that participants in the user study will not hold a white cane while exploring the system. An assistant who is not related to this project should help to ensure the safety of the participants. The problems P0 encountered and the improvements of the prototype are summarized in two aspects: interaction features and flowchart. The resulting flowchart is formulated in Figure 5.
3.3.1. Interaction Features
An acoustic signal recognition module triggers steps like target definition, target confirmation, and option selection in our initial flowchart. Due to language redundancy, P0’s voice input wasn’t always fully recognized or understood. Inspired by call forwarding, we adapted the system to replace most voice commands with button selections, except for target and open question definitions. Consequently, the two buttons on the waist bag, labeled “1” and “2”, handle confirmations and selections: “1” for “yes”, “2” for “no”, and combinations of “1”, “2”, “12”, or “21” for choosing among up to four options. We omitted double-clicking due to the buttons’ limited key travel. We employed a continuous low-frequency beep to indicate the that system is in operation and enhance user’s sense of safety. Additionally, taking cues from Orcam MyEye2 (OrCam Technologies, 2022), we integrated earcons for target detection and input readiness, with beeps at $1Hz$ for system activity and $3Hz$ for model loading and MLLM responses. After confirming targets, users should move their heads to scan and stand still following the earcons. A misalignment issue between the user’s head and body prompted us to add, “Please align your body with your head’s direction.” to each MLLM output to clarify the expected posture.
User ID Gender Age Range Vision Level, Onset Experience of Apps
P0 Male 30-39 Light perception, since about 2004 Be My Eyes, Seeing AI
P1 Female 20-29 Light perception, since about 2022 Be My Eyes, Seeing AI
P2 Male 50-59 Fully blind, since birth Be My Eyes, Seeing AI
P3 Male 20-29 Fully blind, since birth Be My Eyes
P4 Male 20-29 Fully blind, since 2010 Lookout
P5 Female 30-39 Light perception, since birth Seeing AI, Envision
P6 Female 50-59 Fully blind, since about 1989 Seeing AI
P7 Male 70-79 Fully blind, since birth None
P8 Male 30-39 Light perception, since 2015 Seeing AI
Table 1. Participant demographic statistics in Phase II interview with target group.
3.3.2. Flowchart Refinement
Participant P0 noted that route planning (F3) was a unique MLLM feature distinguishing our ObjectFinder from Be My Eyes (Eyes, 2024) and frequently activated this feature. However, remembering the instructions after a single playback proved challenging, so the instructions needed to be repeated. Additionally, P0 might lose orientation after taking several steps toward the target, necessitating additional scene descriptions. To address potential confusion regarding the route, we encourage users to search for the target again. Consequently, we provided options to continue searching, repeat, and scene descriptions in block 6, following the route planning section in block 4. We also introduced open questions in block 6 after scene description (F4) and open question (F5) in block 4 to solicit more detailed feedback. In scenarios where the target object was absent, we added block 5 to inform users of available objects. (see Figure 5)
4. Phase II: Interview Study
During Phase II, we conducted interviews with $8$ participants to understand the technology they use and to identify challenges they face when searching for objects.
4.1. Method
To better understand the target objects, we asked participants about the items they most frequently or laboriously search for in their daily lives, and which objects they would be interested in locating using the system, such as furniture or smaller items. To gauge whether they require inspiration for unexpected targets (C2), we inquired about the objects they would expect to find in an unfamiliar office or living room, and what items they might not recognize unless informed of their existence. We also explored the advantages and disadvantages of the technologies and methods participants currently use for object search or scene understanding. Additionally, we specifically asked whether they could match photos taken by a description-based system, such as Be My Eyes, with the actual environment.
4.1.1. Participants and Procedure
[…]
The interview was conducted at the beginning of the exploratory user study described in Sec. 5. The participants were compensated €$40$ for study participation, and travel costs were reimbursed.
4.1.2. Data Analysis
We analyzed qualitative data using the thematic analysis framework developed by Fereday and Muir-Cochrane (Fereday and Muir-Cochrane, 2006). The first author led the analysis, first reading and re-reading the transcripts, then engaging in coding, and finally organizing the data into themes. The themes were subsequently reviewed by the research team. In this phase, $122$ data points were categorized into $42$ initial codes, which were then distilled and refined into $15$ codes, from which two main themes were crafted.
4.2. Results
Here, we present the two themes crafted from the Phase II interviews, (1) Challenges of Existing Technology for Object Search, and (2) Target Objects and Difficult Cases for Object Search.
4.2.1. Challenges of Existing Technology for Object Search.
The data in Table 1 shows that most participants prefer description-based applications. Notably, only participant P4 reported using the detection-based application Lookout. The use of Be My Eyes for document summarization and volunteer interaction is not included in this table.
Description-based Systems
One challenge for blind people is to take the right photo for a description-based system due to the heavy reliance of photography on the visual domain (C1). One participant commented on Be My Eyes, “if you lose this point, you can receive the wrong truth. A wrong picture you never can be sure, you have to know it if you use these things. ” (P2). Seeing AI and Envision do not help to take the right photo as well, but they do indicate if the photo is of low quality. “There is a Seeing AI app for example. It tells you, ‘No, this corner is not visible’. […] Envision, it doesn’t tell you unless you already take the photo and it processes the photo, then it tells you ‘no it was quite blurry’.” (P5). In terms of learning if the users of Be My Eyes (P0-3) can align the photo with the real environment: all of them answered yes, but participants reported that it is difficult in an unknown environment, e.g., “it’s easier when I know the surroundings, so I can align what’s described. But somewhere I’ve never been before, I find it more difficult.”(P1). Besides, P2 mentioned a practical method to take good photos: “ you can learn to fix: you can go [to] touch the object, and you go one or two steps back, take your camera and focus.”. This suggests that the effectiveness of description-based systems in developing a mental map within an unfamiliar environment without tactile exploration may be limited. In addition, participants were suspicious of AI-generated descriptions (Lee et al., 2022a; Xiao et al., 2024) without knowing “when does it start lying?” (P2). However, the availability of these applications on the cell phone was considered a significant advantage (P1, P3), while privacy concerns must be considered, as everything is “sent to the cloud” (P0).
Other methods for Object Search
As for object-search systems, P2 noted: “at the moment we have no usable object-finding system in most cases.”. However, participants mentioned various methods to search for objects. For example, P4 uses Lookout, but not frequently due to poor internet connection. Some mentioned using tactile exploration and white canes to search for objects in their daily activities (P2, P3, P7). As for other technologies, P2 mentioned adding audio descriptions to QR codes with PenFriend to distinguish private items, such as clothing colors, food containers, or medication. Besides, participants attached AirTags to their keys, although not everything can be equipped with the technology (P2, P6). It enables participants to locate items using sounds. For example, they might use an Apple Watch to ping their phone to find their seat in a beer garden. However, the aforementioned methods do not have the flexibility to define targets flexibly, and cannot inspire users with interesting unexpected discoveries in unfamiliar scenarios (C2).
4.2.2. Target Objects and Difficult Cases for Object Search.
Here, we analyse the role of target objects and difficult cases for object search in both familiar and unfamiliar scenarios.
Familiar Scenarios
In their familiar private spaces, the belongings of participants are well-managed and placed in specific locations. However, search for them can be difficult if objects are moved by other people (P2, P5), or if participants misplace their own items (P5, P8). As P2 mentioned, “ someone was in my office and changed a little thing to put her laptop and something on the table. After that, it took an hour to have my things back.”. P8 also noted, “ if I’m lazy and put it somewhere, then I really know it has to be around here, but it’s not in its usual place.” Participants also reported searching for small objects in their daily lives: More than one participant had previously looked for their smartphone, keys, small electronics (charger, USB stick, etc.), or items such as coffee cups. Sometimes, even though the participants know that objects are nearby, exact location matters, with one participant pointing out that “‘Near’ could be a problem […] your coffee spoiled your laptop.” (P2)
Unfamiliar Scenarios
In unfamiliar environments, participants reported difficulties in knowing what is on top of a table (P3, P5, P8) , such as food, because “you usually don’t feel around whole tabletop.” (P3). In addition, the participants are unlikely to know the paintings and displays hanging on the wall (P0) and the “hidden” objects (P6).
Additionally, knowing what one can use in an unknown scenario can be a challenge (C2). For example, P4 explained that “we once asked ourselves in the student room whether we had air conditioning. […] The question really occupied us for several hours [and] whether it would be warm in the practice room in the summer. […] If we could scan the room ourselves, we could determine ourselves whether climate regulation is present.” In contrast to the typically small target objects in their daily lives, most participants reported looking for large pieces of furniture when they did not have a mental map of an unknown environment , because “I find it good to get oriented quickly, especially in a foreign room.” (P6). P1 typically aimed to find the “next available chair” in an unfamiliar office. Meanwhile, P2 sought to locate the cupboard, bed, and other furniture in an unfamiliar hotel room, noting that plugs were harder to identify. In smaller rooms, participants commented on the value of tactile exploration to investigate their surroundings, with P7 noting that he “go through the room and feel with my hands to see what I can find next.” However, it is important that this is not feasible in all environments. For example, in the specific case of an unfamiliar toilet, “it’s very important to find the objects, because [otherwise] you have to touch everything” (P2).
Functions Passive Information Aquisition Interactive Object Search
Lookout (Google, 2024) Be My Eyes (Eyes, 2024) ObjectFinder_Base ObjectFinder
[…]
Figure 6. Simplified user study procedure focusing on two target objects: a large piece of furniture (an office chair) and a smaller object (a plate of cookies). The user walks into an unknown environment, the office in this example. To sit in front of a desk, he or she should find an office chair first and navigate to it. Then the user sits on the office chair and defines the desk as the region of interest. The user will know what is on the desk through the scene description function, a plate of cookies in this example, and get additional information through open questions.
[…]
This figure illustrates a bird’s-eye view of a room layout and the sequential steps a blind user takes to navigate and interact with the environment using an assistive system. Step 1 involves the user entering the room and commanding the system to ‘Find office chair’. Step 2 initiates a scan of the environment for open-vocabulary object detection. In Step 3, the system provides localization information, indicating ‘office chair found, 2.6 meters, 11 o’clock.’ Step 4 involves route planning with guidance like ‘Six steps forward, follow the edge of the desk, at knee level.’ Step 5 allows the user to sit on the office chair and triggers a scene description: ‘There’s a desk in front of you with a plate of cookies.’ Finally, Step 6 involves prompting the user to ask open questions such as ‘What flavor are the cookies?’, to which the system responds ‘Chocolate!’
5. Phase III: Exploring User Perspectives on ObjectFinder
In Phase III, we conducted an exploratory study with blind people to determine if ObjectFinder can support the object search in unfamiliar environments. In this step of our work, we include our own prototype ObjectFinder, ObjectFinder_Base (a closed-vocabulary baseline prototype), and the commercial systems Be My Eyes (Eyes, 2024) (a description-based application) and Lookout (Google, 2024) (a detection-based application) as points of reference for participant feedback.
5.1. Method
We invited participants P1111-P8888 (Sec. 5.1.1) to explore the two prototypes and the commercial applications in two unknown environments that were set up in our research lab: a space mimicking a living room, and an office. The functional differences between the tested systems are listed in Table 2. ObjectFinder_Base differs from ObjectFinder that it only recognizes a limited number of objects with fixed wording. The systems are described more detailed in Sec. 5.1.2.
In order to assess users’ experiences with the systems and to find out whether an open-vocabulary offers an advantage over a closed-vocabulary approach, we compared ObjectFinder with ObjectFinder_Base in terms of: expectations regarding the functionality, mental map that users formed of the environment after using one system, cognitive load and satisfaction. We used semi-structured interviews featuring Likert-scales for function evaluation and open ended questions. Questions ranged from “In which areas do you believe our system could be improved?” and “Do you want more or less information?”. Additionally, we included the NASA-TLX (Hart, 2006) for assessing the cognitive load to accomplish the tasks, and the Net Promoter Score (NPS) (Reichheld, 2003), a measure that reflects the likelihood if the users would like to introduce the system to their community. The exploration of the commercial applications was conducted separately, followed by a short semi-structured interview and the Net Promoter Score (Reichheld, 2003).
5.1.1. Participants and Procedure
The participants were the same eight people who took part in the interview described in Phase II (see Sec.  4.1.1). For scene understanding, six of the participants had previous experience using description-based applications such as Seeing AI, Be My Eyes, and Envision, while only one participant uses a detection-based application, Lookout. In Table 1, we consider only the scene description or exploration feature of these applications, while usage of the applications for other purposes, such as reading documents, is not included.
[…]
Participants were compensated with €$40$ for study participation, and reimbursed for the travel cost.
5.1.2. Apparatus
In this section, we describe the systems we used in the evaluation: ObjectFinder, ObjectFinder_Base, Be My Eyes and Lookout.
ObjectFinder and ObjectFinder_Base. The hardware and flowchart of ObjectFinder and ObjectFinder_Base are identical, except that the baseline only detects object categories in COCO2017 (Lin et al., 2014). For ObjectFinder_Base, we set the confidence level threshold to $0.7$ to filter out falsely detected objects. In contrast, the YOLO-World used in ObjectFinder can detect any target outside the training dataset, resulting in potentially lower confidence levels. We set the confidence threshold for YOLO-World to $0.3$ and tested it with $20$ objects, including those used in our user studies. It detected all targets without any false positives, even when the objects were located $8$ meters away. Since some target objects are undetectable by the baseline system, we added additional options between blocks 1 and 2, as shown on the right side of Figure 5, to allow participants to trigger scene descriptions multiple times and determine if the target objects are present in the region in front of them.
Lookout. The Explore mode of Lookout, which was tested in this study, detects the nearest object in front of the camera and provides real-time outputs including object categories, quantities, and associated text.
Be My Eyes. Be My Eyes integrates GPT-4 into its Be My AI feature for a vivid scene description, replacing the need for sighted volunteers.
Lookout and Be My Eyes reflect the two main parts of our prototype: object detection and scene description. However, users cannot actively select their targets, obtain localization information and navigation instructions with these two applications, but instead passively receive information chosen by the system.
5.1.3. Scenario Exploration
In each scenario (living room or office), participants were asked to find six target objects: three large pieces of furniture, in order to build a mental map of the unfamiliar environment, followed by three smaller objects found on the desk or coffee table. Figure 6 describes a simplified procedure involving the five functions F1-F5. Table 3 specifies the target objects that need to be found in sequence and the initial MLLM functions to be triggered when these target objects are detected. The layout and the order of targets for search are shown in Figure 7(b). We are the first to engineer prompts that generate route planning instructions for blind individuals, guiding them to objects. Therefore, our primary focus is on testing the route planning function (F3). Since the closed-vocabulary ObjectFinder_Base can only detect a limited number of target objects (Lin et al., 2014), we categorize the six target objects in each scenario into three groups: unrelated to (two objects), related to (three objects), and exact in (one object) COCO2017. As for the related targets, we will hint to the participants to look for the related objects in COCO2017 when they are using the baseline, but they will experience a vocabulary gap between the target definition and MLLM information. As for the unrelated objects, the participant cannot even define the target objects (C1). So we provide the option that the participant can trigger the scene description function several times to find the objects, as illustrated on the right side of Figure 5. Cookies are the bonus target, and we observe whether the participants can recognize it themselves to validate the capability of our system to inspire the detection of unexpected targets (C2). Figure 8 illustrates several examples of how the participants detected the target objects and received the outputs of the system during the user study.
The commercial applications, Lookout and Be My Eyes, which were already familiar to the participants and not specifically designed for object search tasks, were briefly used in the living room scenario to locate a fan and a teapot as a way to remind participants of their functions.
(a) Living Room.
(b) Office.
Figure 7. Layout of the two scenarios and the order of target objects. The 7th object in the office, a plate of cookies, is the bonus, to determine if participants are aware of its existence through the system. The photos are taken at the starting points of the task in each scenario.
[…]
This figure provides a visual representation of the layout for two scenarios: a living room and an office. It details the sequence and locations of various target objects within these settings.
Scenarios Target Objects Names in COCO (Lin et al., 2014) dataset
Office Furnitures trash binF3, deskF4, office chairF3 ¡None¿, dining table, chair
Smaller objects monitorF5, keyboardF3, headphoneF3 (cookiesF3) TV, keyboard, ¡None¿ (¡None¿)
Living Room Furnitures fanF3, coffee tableF4, sofaF3 ¡None¿, dining table, couch
Smaller objects teapotF3, bananaF3, flowerF5 ¡None¿, banana, potted plant
Table 3. Target objects in each scenario. The superscript objectF denotes the function that is activated first upon finding the target objects. F3: route planning; F4: scene description; F5: open questions.
Figure 8. Examples of functions F3-F5 include triggering route planning (F3) to reach both a large piece of furniture and a small item, using scene description (F4) to describe the coffee table, and employing open questions (F5) to inquire about details. Postures (first column) and key frames (second column) are visualized, with audio input and output also provided.
[…]
This figure presents four real examples from a user study illustrating how participants navigate to and interact with both large pieces of furniture and small objects, as well as how they describe and inquire about items in a room setting. Example 1 involves navigating to a fan (5.5 meters, 11 o’clock) with detailed route planning instructions: ’Align your body in the direction of your head. Move straight ahead for about three steps. Then move slightly diagonally to the right for about two more steps. Then fan is at floor level and is about one meter in front of you on the right side.’ Example 2 describes finding headphones detected at (0.7 meters, 1 o’clock) with instructions: ’Align your body with the direction your head is facing. Reach out directly in front of you, slightly to the right, to find the edge of the desk. Once you find the edge, move your hand to the right for about 6 to 8 inches to touch the plate with cookies. The headphones are located just to the right of this plate. You will feel them resting on the desk.’ Example 3 details the items on and around a coffee table (2.4 meters, 11 o’clock): ’A sofa with cushions is positioned to the left of the coffee table. On the coffee table, there are bananas, a cup, and a remote control. A plant is situated behind and to the right of the sofa.’ Example 4 shows an inquiry about a flower (1.0 meters, 10 o’clock): Participants ask ’What colour is the flower?’ and ObjectFinder responds ’The flower has petals that are a mix of pink and yellow. The pink is more dominant on the edges of the petals, while the yellow is situated toward the center.’
5.1.4. Data Analysis
In this phase, we have both qualitative data and quantitative data. For qualitative data, the user study transcripts were analyzed using the hybrid process of inductive and deductive thematic analysis proposed by Fereday and Muir-Cochrane (Fereday and Muir-Cochrane, 2006). Analysis was led by the first author who read the transcript repeatedly to achieve familiarisation, followed by multiple rounds of coding. Beyond data-driven inductive coding, we also applied deductive coding where it yielded meaningful insights into the system’s capacity to identify regions of interest (C1) and to facilitate the discovery of unexpected targets (C2). In a workshop, the research team assigned $243$ data points to $69$ codes, which were further refined to $12$ codes, and finally, four themes were crafted and will be presented in Sec. 5.2.
Due to the small size of the user group, inferential statistical analysis of the quantitative data is not meaningful. Therefore, we limit our comparison to descriptive statistics.
5.2. Results
Here, we present the following four key themes crafted from the interview findings, which focus on the users’ Information Needs for Object Search, their experiences with Mental Mapping, their System Perception and Preferences as well as System Requirements.
5.2.1. Information Needs for Object Search
In this section, we analyse the amount of information the participants preferred, and the obligatory and optional information for object search.
Amount of Information
Regarding the amount of information provided, three participants found it adequate for their needs. However, when asked about their preference for more or less information, responses were divided: half preferred more, while the other half favored less. We noted that preferences for information quantity relate to individual processing styles: while some participants could ignore excess information, others felt overwhelmed. The participants chose to receive more information, generally expressing a preference for as much as possible. As one participant noted, “as much as you can get. It’s completely blank for me, so the more I have, the better.” (P5). P4 expressed that “things that don’t interest me, I can just ignore.” Conversely, P6 and P7 explained their preference for less information, attributing it to not being accustomed to processing such a large amount of visual information and feeling overwhelmed by it. P8 suggested that the information provided could perhaps be reduced after getting the first overview: “it could maybe be a little bit less, so if you know once there are some objects on the table, then you don’t really need this information the second or third time unless you ask the system what’s on the table.” Additionally, we observed that the amount of information varies across functions. Be My Eyes, used solely for scene description, received praise from three participants for its “detailed” information, although one deemed it excessive. In contrast, participants noted that our ObjectFinder delivered more information than necessary for efficient navigation.
Obligatory Information
Participants (P6-8) highlighted the importance of localization information (distance and direction) after scene exploration. The interpretation of the distance should be clear and intuitive. P1 highlighted the usefulness of different distance measures noting, “I think the steps are good for like when it’s really near, so it’s just a few steps. But if there are longer distances. I think sometimes meters might be more useful […] So it sometimes has like $0.1$ meter […] you could have said ‘right in front of you’ or ‘one step’ like that.” Participants mentioned that the distances reported were inaccurate, appearing larger than they perceived. Upon reviewing the video, we observed that some discrepancies were caused by measurements taken from the head to the object, rather than horizontally, as shown in Figure 11. This inaccuracy may be more pronounced in our small indoor settings. To describe the direction, P5 suggested to “specify a $20$ centimeters margin”, noting that if an object is $10$ centimeters to the left or right, it’s still considered in the front.
Contextual information around the target is crucial for object search, such as relationships among surrounding objects and their location information. For example, P3 noted, “it’s good that the objects around it are announced, so you have some idea where the object you want to find is in relation to other objects.” A description should also include the distance which says something about the user as a reference point to the object. P8 suggested to improve ObjectFinder by incorporating distance into the MLLM output, “ when you make the description, it holds everything but it does not really talk about the distance.”
Optional Information
Participants expressed mixed feelings about the relevance of color and alert information, which emphasizes the importance of personalization. For example, P5 and P6 appreciated the unsolicited color details. “It told me the color of the remote control without my asking […] it helps me to visualize the environment around.” (P5). In contrast, P2 criticized the excessive information, remarking, “I want to walk from A to B, and I don’t want a literature presentation of the color and the landscape description which can fill books.” Regarding the alert information, P2 valued the clarity of certain descriptions: “what I liked was the description. It directly tells people to be careful ‘move your arms’, and I think it was a very clear description.” However, P3 found it superfluous, commenting, “about the whole flavor text about ‘not bumping into an object’, ‘sitting down’, ‘carefully turning the chair toward me first’. This is all not really necessary.”
Figure 9. Evaluation of the five functions in terms of importance and level of interest, using Likert-scale scores, before and after exploring two scenarios. F1: object detection; F2: localization information; F3: route planning; F4: scene description; F5: open questions.
[…]
This bar chart compares Likert-scale scores for five functions before and after an exploratory user study, assessing importance and level of interest. Most scores increased post-study, except for object detection, which slightly decreased from 4.75 to 4.5. Localization information received the highest scores, with importance rising from 4.38 to 4.88 and interest from 3.88 to 4.75. Scores for the first three functions exceeded 4, while the remaining two stayed below 4.
Figure 10. Net Promoter Scores (NPS) for different systems, assessed by participants P1-P8 while exploring scenarios.
[…]
The bar chart of the net promoter scores for different systems: ObjectFinder (M=6.75, SD=2.66), ObjectFinder_Base (M= 5.88, SD=2.42), ObjectFinder with efficient hardware (M=8.13, SD=2.01), Be My Eyes (M=7.25, SD=3.01), Lookout (M=5.31, SD=2.43).
Figure 11. Special cases. Upper half: Initially, the participant did not detect the coffee table as an obstacle to the sofa. Subsequently, the participant stepped backward to detect the coffee table instead of lowering the head. Bottom Left: The participant needed cues to lower the head to detect items on the desk. Bottom Right: The participant felt uneasy about the distance ($1.5$ meters) to the office chair because the measurement was taken from the head to the chair, not horizontally.
[…]
This visualization depicts the postures and key frames for special cases where participants needed cues to lower their heads, and felt uncertain about the distance. In the example shown in the lower right part of the figure, a participant who is about 1.7 meters tall detects a chair in her vicinity (0.5 meters away), yet the distance is calculated to be 1.5 meters from the glasses to the chair.
5.2.2. Mental Mapping
There is some evidence that participants obtained more detailed environmental information through the use of ObjectFinder. P1 noted:“I think it was very good [to know] where I am. When it said where my target object was, it also told me the surrounding objects, because maybe if I find the surrounding object first, then I know, OK I’m close. ” P6 noted the benefits of discovering other usable items (C2) and understanding their arrangement, “I didn’t know there was another chair there. It gives more information about objects and their arrangement.” Participants wished to get an overview of unfamiliar rooms without receiving too many details. P8 explained, “I could imagine if you really don’t know the room, and you want to first [have an] overview what is in the room, it was pretty detailed. ”
ObjectFinder’s ability to locate regions of interest is a key feature that facilitates detailed descriptions, aiding in navigation and discovery without physical search (C1, C2). P6 particularly valued this feature, noting: “if it’s a new desk, I don’t have to feel around to know where the computer is. I don’t have to move far right or left to find the position easily.” Similarly, P1 found the system’s detailed output helpful for discovery (C2), “When I first got into the scene, I got a very detailed description of the coffee table […] so I knew what I could expect to find there.” Moreover, the wearable design of ObjectFinder with glasses that capture egocentric images significantly improves participants’ sense of direction (C1). “I found it easier to think about which direction.” (P6).
This highlights the potential of ObjectFinder to support users in building a mental map of an unknown space. This was more difficult using the existing commercial systems. For example, after using Be My AI, P8 mentioned, ”Be My AI had no information regarding the distances, so you have […] a less sense of location ”. As previously analyzed, localization information (distance and direction) and the relationships between surrounding objects are crucial for effective object search. The lack of localization information (P2, P8) and the absence of context regarding object relationships (P2) were mentioned as reasons why participants did not prefer Lookout, highlighting the potential for solutions that can combine multiple aspects of object search.
5.2.3. System Perception and Preferences
In this section, we analyze the general perception of ObjectFinder, the user preference with the systems, and the cognitive load to finish the tasks.
General Perception of ObjectFinder
Generally, five out of eight participants expressed excitement about ObjectFinder after exploring it. “I was very positively surprised at how good it works, and how easy it is to get information and to find myself in the scene.” (P1). ObjectFinder was also appreciated for accurate object searches without physical touch. Subsequently, participants rated the system’s helpfulness ($\mu=4.13,\sigma=0.641$) and their independence ($\mu=4.31,\sigma=1.01$) with ObjectFinder in an unfamiliar environment, with both ratings based on $5$-point Likert scales and averaged over Through ObjectFinder, participants discovered more than just target items, finding inspiration in unexpected objects, both small and large (C2). For example, “I’m sort of inspired because now I know what options I have, instead of walking in and just looking for my phone because I know it must be there, but I don’t know what else could be there […] I was inspired to look for, for example, bananas and stuff like that.”. (P1). However, P3 noted that searching with ObjectFinder might be slower than tactile exploration in our compact indoor setting, noted, “there are so many steps putting it up and searching for the thing, and maybe it doesn’t even find it (ObjectFinder_Base).” The search could be accelerated by using ObjectFinder alongside the cane, taking advantage of the cane’s large radius (P2).
User Preferences and Cognitive Load
The NPS for the prototypes does not vary substantially between the scenarios, office ($\mu=6.25,\sigma=2.49$) and living room ($\mu=6.38,\sigma=2.67$), with scores being generally high, and not lagging behind those allocated to commercially available systems (see Figure 10). With respect to differences between the systems and implications for user preferences, participants appreciated that they could actively define the target objects with ObjectFinder, making it more reliable in searching for specific items. As P6 mentioned, “what I also liked about this system is speaking to it, I find that more targeted.”. P3 echoed this sentiment, highlighting the biggest advantage: “The biggest advantage […] contrary to Be My AI is that you can specify what you want to find.” (C1).
As shown in Table 4, the study also examined cognitive load, showing that it was relatively low across system types and scenarios. Retrieving keyframes for MLLM information and participant feedback reveals that the landscape orientation of the camera on the glasses, contrasting with participants’ usual practice of taking portrait photos using commercial applications, and the presence of relatively low furniture in the living room often result in the prototype failing to recognize the coffee table as an obstacle to the sofa, as illustrated in Figure 11. The exploratory experience with the prototype was not comprehensive enough. P7 mentioned that F4 and F5 are insufficiently tested in our study, and added: “one needs to get used to it, it requires some practice, and then the search will be very precise.” P5 mentioned that she was not accustomed to the earcon indicating to stand still.
Sub-scale Systems Scenarios
ObjectFinder ObjectFinder_Base Office Living Room
Mean Std Dev Mean Std Dev Mean Std Dev Mean Std Dev
Mental Demand 4.25 2.96 5.38 3.42 4.50 2.93 5.13 3.52
Physical Demand 2.44 1.76 3.00 2.39 2.69 2.25 2.75 1.98
Temporal 2.63 2.00 2.63 1.30 2.13 1.46 3.13 1.73
Performance 4.69 2.02 7.25 2.38 4.94 2.43 7.00 2.27
Effort 4.13 2.03 5.38 3.11 4.75 3.11 4.75 2.25
Frustration 4.13 3.48 4.38 3.07 4.00 3.46 4.50 3.07
Table 4. Assessment of workload required to complete tasks with two systems (ObjectFinder vs. ObjectFinder_Base) and in two scenarios (Office vs. Living Room) using NASA-TLX. The scores range from 1 (very low) to 21 (very high).
5.2.4. System Requirements
In this section, we analyze the system requirements, i.e. interaction features, software and hardware requirements.
Interaction Features.
It seems that interaction features are assessed differently among participants. P4 felt pressing buttons was faster than using a touchscreen, though half of the participants found the button codes for option selection unusual. Suggestions included maintaining a unified button combination for each function (P3) or using “four different pressings” (P8). P1 preferred pressing the button without waiting for the entire list to be read aloud.
Additionally, P8 described the voice commands as “pretty, pretty easy”, while P4 expressed concerns with programs that solely accept voice input, suggesting “maybe it’s better to have the ability to switch between a list of maybe recognized objects and voice commands.” Besides the current earcons, P6 suggested creating an additional earcon for option confirmation, rather than using spoken text.
As mentioned before, some participants preferred more information, while they suggested “implementing a skip option” (P1) and the ability to “switch information on and off optionally” (P2). P4 suggested adding a main menu to easily switch between object search (detection and navigation) and scene description. The interaction feature of Lookout was highlighted by P1: “I don’t have to take a picture and wait for a response.”
Software Requirements
We asked participants to rate the functions before and after using ObjectFinder and ObjectFinder_Base in the explored scenarios, in terms of their perceived importance and level of interest, as captured by Likert scores (Figure 9). From the ratings, the functions object detection, localization information and route planning are important for the object search task ($\mu{>}4.0$), accompanied by a reduced standard error. Among the five functions evaluated, localization information was rated as the most important.
Compared with the commercial applications, P8 mentioned the advantage of ObjectFinder is “the distance and also the guiding function which is not really available for the other both apps.” (P8). We note that the software should be fast and make fewer mistakes. P2 mentioned that receiving explanations from a sighted person on Be My Eyes “works better because it’s without delay.” Half of the participants mentioned frequent misidentifications of Lookout was not preferred. As P5 noted, “it misidentified objects […], like there’s no dishwasher,” and at times, the teapot was categorized as a helmet or mouse.
Hardware Requirements
Five participants mentioned that the hardware was heavy for long-term use. We asked participants to imagine that the ObjectFinder prototype was implemented on more efficient hardware (Waisberg et al., 2024), such as the Ray-Ban glasses for image capturing paired with an iPhone for data processing, while actually trying on a pair of Ray-Ban glasses to evaluate their feel. With these enhancements, the Net Promoter Score (NPS) of ObjectFinder increased from ${\mu}{=}6.750,{\sigma}{=}2.659$ to ${\mu}{=}8.125,{\sigma}{=}2.013$, as illustrated in Figure 10. In comparison, a common advantage of the applications, Be My AI and Lookout, is the availability on the smartphone.
Participants noted the distinct advantages and disadvantages of capturing scenarios with glasses compared to a cellphone. Three participants experienced reduced mental effort in determining the camera’s orientation when using glasses that capture the egocentric view. “It’s always harder for me to think exactly about what the phone is capturing with the phone camera, I found that better with the glasses.” (P6). On the contrary, users required external cues to detect objects in the lower part of their viewing field. “The possibility of overlooking some obstacles like the coffee table […] if I hadn’t known that there was the table, I would have just run into it, which is not that nice.” (P1). To detect the obstacles, such as coffee table, participants should lower their heads rather than step backward. Two participants, who explored the office using ObjectFinder_Base, which couldn’t locate the region of interest for certain objects, without lowering their heads, failed to detect items on the desk and remained unaware of a plate of cookies also present there. (C2) Additionally, using glasses poses a challenge since they represent an additional item to carry alongside smartphones. Consequently, a pair of glasses is another item for them to carry and potentially forget, alongside their smartphones (P4, P5). The price of the glasses was also mentioned to be considered.
6. Discussion
In this section, we summarize our key findings to address our research questions, discuss potential use cases for the object-search system, and reflect on requirements to be addressed by future work for object-search systems to be of use to people who are blind.
RQ1: How can an open-vocabulary prototype integrate detection-based and description-based technologies to support object search for blind people, and what challenges are associated with its design and development?
Overall, our experience with the design and implementation of ObjectFinder suggests that it is feasible to create assistive technology for object search that combines detection-based and description-based technologies, and that can be leveraged by blind people wishing to examine unfamiliar environments. In particular, our work shows that such a system can facilitate exploration of different settings (i.e., living and office environments), and that it can be a first step toward building a mental map of an unknown space. In particular, the description part of the ObjectFinder prototype offered an overview that helped participants understand the environment, know what to expect in the scenario, and discover unknown objects for subsequent searches. Compared to the description-only and detection-only applications, systems with the design of ObjectFinder have potential to provide a guiding function and the comprehensive information necessary for object search: localization information (distance and direction) and the relationship among the surrounding objects. However, our findings also show that there are several design challenges that remain unaddressed. For example, open questions regarding hardware design and interaction features remain, and addressing these is crucial when creating technology that people are interested in integrating into their daily lives. With respect to the information provided by the system, results suggest that the amount of information and whether to make additional details optional for object search, such as color and alert information, needs to be fine-tuned based on participants’ information processing strategies.
RQ2: What are users’ experiences and preferences regarding the open-vocabulary interactive object-search system, specifically comparing the description-based and the detection-based, with the hybrid ObjectFinder prototype?
Given the absence of existing technology to assist blind individuals in object search, most participants expressed enthusiasm for this open-vocabulary interactive object-search prototype. Participant ratings reflect satisfaction with the prototypical integration of functions within ObjectFinder, and suggest that cognitive load during system use was acceptable. In comparison with exclusively description-based or detection-based systems, our prototype effectively defines the target object, locates regions of interest, and delivers essential localization information and route planning, with interview results showing that this was appreciated by users. With respect to user experience with specific features of ObjectFinder, our work shows that detection of small objects was well-supported, whetheras the user experience when exploring larger space was less smooth, with open questions regarding user safety remaining (i.e., better obstacle recognition).
6.1. Future Directions for the Development of Object-Search Systems
Our research highlights that object-search systems have potential as an assistive technology for people who are blind. Here, we want to highlight future directions for the development of such systems, outlining challenges and opportunity for the Human-Computer Interaction and accessibility research communities.
6.1.1. Potential Features for Future Integration
In our study, we uncovered a range of features and characteristics of ObjectFinder that offer interesting avenues for future work, either through iteration on our system, or integration in other, comparable systems. Here, we give an overview of the most relevant aspects.
Providing better and more tailored descriptions. Providing descriptive overviews of unfamiliar environments is essential for revealing unexpected objects and should be relative to the user’s position. Based on our findings, we recommend to include both distance and direction in descriptions, with far objects quantified in meters and nearer objects described in steps or as ‘right in front of you’. These details can facilitate room exploration, help in building mental maps, and assist in orientation.
Understanding advantages and drawbacks of additional information. Although the essential information for object search, localization information and relationships among surrounding objects, is well-defined, participants had mixed feelings about additional details like color and alert information. Consequently, the future object-search system should allow users to select the amount of information by incorporating options to skip information or to stop and continue information output. Likewise, future work should explore user preferences for the types of information to be included, and implications of additional unrequested descriptions for user experience and aspects such as cognitive load.
Improving system reliability and integration with other assistive technology. Unsurprisingly, participants expressed a preference for a system that is error-free and can accurately locate searched objects. From a system requirements perspective, the camera should capture both nearby objects at a lower angle and objects at a distance. The system should be lightweight and offer intuitive interaction features. For our prototype, participants appreciated the use of a button for option selection over a touchscreen, as well as voice commands. Additionally, earcons should be intuitive, while a training session to acquaint users with the meanings of earcons at the outset is advisable. The design of the system should serve as an extension to the cane, taking advantage of its large radius.
Addressing portability and social accessibility of the system. Our results show that participants reflected on the hardware included in the current iteration of the prototype, which is clearly visible, and takes up significant space when attached to the user’s body. Here, participants expressed a preference for a smartphone-based solution. On the one hand, this is a sustainable approach that leverages hardware already in the possession of users. On the other hand, this may address concerns with respect to social accessibility, i.e., the visibility of assistive technology to others, and associated stigma (Shinohara and Wobbrock, 2016).
6.1.2. Tensions and Concerns Regarding Vision- and AI-based Assistive Technology
There are tensions and concerns that need to be resolved for systems such as ObjectFinder effectively and safely supporting object finding.
User habits and needs regarding lighting conditions for camera-based systems. With respect to technical requirements of current approaches to computer vision systems, we note that lighting conditions are crucial for camera-based systems. However, this conflicts with the fact that light plays a different role in the lives of blind people: Legally blind persons may not use light in their homes and workplaces in the same way as (typically) sighted system developers would anticipate  (Gonzalez Penuela et al., 2024), and people who do have residual vision may not find lighting conditions required by camera systems comfortable.
Addressing safety concerns in the context of AI. Likewise, we noted instances in which ObjectFinder did not recognize furniture, e.g., when living room furnishings were lower than those in office environments, and not in view of the camera worn by participants. While there are established strategies for people who are blind or have low vision to use gaze for environmental scanning (Riddering, 2023), no specific scanning strategies for smart glasses tailored to blind individuals currently exist. With advancements in wearable systems and smart glasses, researching and integrating these specific scanning techniques into mobility training is both viable and beneficial. Moreover, incorporating sensors and cameras with wider angles, like LiDAR (Liu et al., 2021a) and omnidirectional camera (Kawaharazuka et al., 2024), could enhance scene perception over larger areas, reducing the need for physical scanning efforts. However, despite potential technological solutions (which may come with new challenges), this specific instance highlights tensions around safety: On the one hand, users are invited to rely on systems for object detection, on the other hand, it is known that vision- and AI-based systems can be unreliable in specific situations (e.g., in the context of autonomous driving (Cummings and Bauchwitz, 2024)), and even more so in the context of disability (e.g., see (Kane et al., 2020)). Thus, there remains a tension between what AI-based assistive technology seeks to offer, and what it can realistically provide, which is an aspect that needs to be communicated with nuance, and should be negotiated together with target audiences in the context of future work.
Understanding the limitations of technology for object search. Finally, the Human-Computer Interaction and accessibility communities have previously discussed issues surrounding technologies focusing on independence of users. In particular, Vincenzi (Vincenzi, 2021) contributed a critical appraisal of assistive technology for navigation of blind people, suggesting that there were instances where working with other people was more relevant than applying a technology-based solution. In this context, the principle of interdependence (Bennett et al., 2018) is relevant, i.e., the fact that we all exist within relationship with our environment, and that assistive technology should not only consider the individual user but also the (social) context within which it exists, and implications of its design under consideration of opportunities to create collective access. Here, we need to ask critical questions around specific features of ObjectFinder, and we want to leave you with one example: Is it really necessary for a blind person to use the system in the workplace, or could non-disabled colleagues make a bigger effort to not misplace or alter their desk?
7. Limitations and Future Work
There are a few limitations that need to be taken into account when interpreting our findings. This research, conducted in small indoor settings of our lab, does not investigate the effectiveness of ObjectFinder in larger, real-world public, private, or outdoor environments. We therefore reached only participants from nearby regions. The requirement for a WiFi connection by ObjectFinder, due to its use of GPT-4, limits its operational range and may introduce security concerns akin to those observed with Be My Eyes. Although ObjectFinder possesses MLLM capabilities to infer surroundings beyond the immediate view, it lacks the ability to precisely remember explored spaces, rendering every space unfamiliar to its current version and thus providing known information to the user. Additionally, while participants preferred our ObjectFinder when paired with efficient hardware (Waisberg et al., 2024), we only had them try the hardware and imagine its implementation, as it was not actually deployed on more efficient hardware.
We see several avenues for future work. With the development of smart glasses, we will pursue collaborations with companies to create a more comfortable and affordable system, which will integrate ObjectFinder’s functionality to support object search and build up mental maps. After addressing the challenges and implementing the suggestions from participants, we plan to conduct a larger-scale quantitative evaluation of the system with blind individuals in real private and public scenarios across different cities. This approach would provide an opportunity to engage a larger number of participants and to evaluate the effectiveness of ObjectFinder in real-world scenarios, which we initially opted not to do at scale. Our first goal was to assess whether the open-vocabulary interactive object-search system addresses existing challenges in daily life and whether our implementation delivers an independent and confident experience to the target group in controlled settings.
8. Conclusion
In this work, we explored the design and development of a prototype that combines detection and description to enable open-vocabulary interactive object search for blind people. With our prototype, we address shortcomings of existing systems that are either description- or detection-based: locating regions of interest and discovering incidental targets. Results of the exploratory user study suggest that the approach is promising, offering blind people a tool to develop a better understanding of unfamiliar spaces. Overall, our work represents an initial step toward developing AI-based assistive technology that supports object search, providing first insights into user requirements and application challenges. Here, we hope that our work will encourage and facilitate further development of object-search systems, and that it will inspire future studies into the experiences that blind people have with such technologies.
9. Acknowledgements
This work was supported in part by the Ministry of Science, Research and the Arts of Baden-Wurttemberg (MWK) through the Cooperative Graduate School Accessibility through AI-based Assistive Technology (KATE) under Grant BW6-03, in part by funding from the pilot program Core-Informatics of the Helmholtz Association (HGF), and in part by Karlsruhe House of Young Scientists (KHYS). We thank HoreKA@KIT, HAICORE@KIT, and bwHPC supercomputer partitions. We adjusted in some cases the grammar of the paper using ChatGPT.

### v2

@@
-ObjectFinder: Open-Vocabulary Assistive System for Interactive Object Search by Blind People
+ObjectFinder: An Open-Vocabulary Assistive System for Interactive Object Search by Blind People
 Abstract.
-Assistive technology can be leveraged by blind people when searching for objects in their daily lives.
-We created ObjectFinder, an open-vocabulary interactive object-search prototype, which combines object detection with scene description and navigation.
-It enables blind persons to detect and navigate to objects of their choice.
-Our approach used co-design for the development of the prototype.
-We further conducted need-finding interviews to better understand challenges in object search, followed by a study with the ObjectFinder prototype in a laboratory setting simulating a living room and an office, with eight blind users.
-Additionally, we compared the prototype with BeMyEyes and Lookout for object search.
-We found that most participants felt more independent with ObjectFinder and preferred it over the baselines when deployed on more efficient hardware, as it enhances mental mapping and allows for active target definition.
-Moreover, we identified factors for future directions for the development of object-search systems.
-People with visual impairments, Assistive technology,
-copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Accessibility systems and toolsccs: Computing methodologies Computer vision tasks
+Searching for objects in unfamiliar scenarios is a challenging task for blind people.
+It involves specifying the target object, detecting it, and then gathering detailed information according to the user’s intent.
+However, existing description- and detection-based assistive technologies do not sufficiently support the multifaceted nature of interactive object search tasks.
+We present ObjectFinder, an open-vocabulary wearable assistive system for interactive object search by blind people.
+ObjectFinder allows users to query target objects using flexible wording.
+Once the target object is detected, it provides egocentric localization information in real-time, including distance and direction.
+Users can then initiate different branches to gather detailed information based on their intent towards the target object, such as navigating to it or perceiving its surroundings.
+ObjectFinder is powered by a seamless combination of open-vocabulary models, namely an open-vocabulary object detector and a multimodal large language model.
+The ObjectFinder design concept and its development were carried out in collaboration with a blind co-designer.
+To evaluate ObjectFinder, we conducted an exploratory user study with eight blind participants.
+We compared ObjectFinder to BeMyAI and Google Lookout, popular description- and detection-based assistive applications.
+Our findings indicate that most participants felt more independent with ObjectFinder and preferred it for object search, as it enhanced scene context gathering and navigation, and allowed for active target identification.
+Finally, we discuss the implications for future assistive systems to support interactive object search.
+copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
 1. Introduction
-We all regularly visit unfamiliar buildings and shared spaces, and need to deal with uncertainty about our environment.
-This also includes blind people  (Müller et al., 2022; Zhang et al., 2021), for whom rapidly changing environments can be challenging to navigate, e.g., finding the way around a room, detecting unexpected obstacles (Alkhanifer and Ludi, 2015; Jeamwatthanachai et al., 2019) and searching for objects without that may have been moved by others.
-When first entering a new or changed environment, it is important to get an overview of the arrangement of furniture, sanitary facilities, and room installations in order to build up a mental map of a room (Façanha et al., 2020).
-Here, commonly leveraged haptic exploration of the environment can have safety implications due to the need for physical contact with potentially hazardous objects or surfaces (Zheng et al., 2024; Gonzalez Penuela et al., 2024), such as knifes and hot drinks.
-Assistive technology has attempted to fill this gap, with some applications (Eyes, 2024; Aira Company, 2024) connecting blind individuals with remote volunteers, enforcing reliance on non-disabled persons for assistance.
-Other approaches leverage AI  (Eyes, 2024; Achiam et al., 2023; Corporation, 2024; Lee et al., 2022a; Liu et al., 2023) to facilitate exploration of unfamiliar environments.
-Such technologies can be divided into two categories: description-based and detection-based systems.
-Description-based systems provide either a vivid and detailed descriptions of a photo (Eyes, 2024) or brief captions (Corporation, 2024; Lee et al., 2022a; Liu et al., 2023).
-However, these systems are not able to localize a specific object in an unfamiliar environment while ensuring that the target object is in the image area, as illustrated on the left side of Figure 1 (Challenge 1, C1).
-As far as detection-based systems are concerned (Google, 2024; V7, 2024; Zheng et al., 2024), current technological approaches only allow the search for specific objects (Constantinescu et al., 2020a; Yi et al., 2013; Schauerte et al., 2012; Ahmetovic et al., 2020; Kacorri et al., 2017; Google, 2024; V7, 2024).
-Therefore, when using current detection-based systems in an unfamiliar environment, blind users may not know what there is in a room and miss unexpected items that could be of interest, such as a plate of cookies in an unknown office, as shown on the right side of Figure 1 (Challenge 2, C2).
-To address these challenges, we designed ObjectFinder, an interactive prototype for open-vocabulary object search by combining both description-based and detection-based methods.
-The integration of Open-Vocabulary Object Detectors (OVOD) (Cheng et al., 2024; Li et al., 2024) and Multimodal Large Language Models (MLLM) (Achiam et al., 2023; Liu et al., 2024) in ObjectFinder makes it possible to develop an interactive process ranging from object detection to description: input any target to get detailed information about the surrounding region of interest.
-This, in turn, identifies potentially interesting and unexpected targets, which can then be searched for in subsequent loops.
-Through our work, we seek to answer the following two research questions:
-  • RQ1: How can an open-vocabulary prototype integrate detection-based and description-based technologies to support object search for blind people, and what challenges are associated with its design and development?
-  • RQ2: What are users’ experience and preferences regarding the open-vocabulary interactive object search system, specifically comparing the description-based and the detection-based, with the combined (ObjectFinder) systems?
-We structured our work into three phases.
-In the first phase, we co-designed with a blind person throughout the conception and development of our system and study design.
-Thus, we first identified a use case that includes the following functions: target object detection (F1), localization (F2), route planning (F3), scene description (F4), and open questions (F5).
-Since existing description-based and detection-based systems are not tailored for object-search tasks, we developed a competitive baseline, ObjectFinder_Base, with limited detectable objects.
-Then, we conducted four half-hour user-in-the-loop sessions with our blind co-designer, inviting him to explore features in different scenarios (a living room and an office), leading to the refinement of ObjectFinder.
-In the second phase, we conducted a semi-structured interview with eight blind participants to understand their daily targets of object search with their currently used methods, and challenges they encounter in unfamiliar or changing environments.
-In the third phase, we conducted an exploratory evaluation of ObjectFinder inviting the same eight participants to engage with the system, and take part in a semi-structured interview afterwards.
-Through thematic analysis (Fereday and Muir-Cochrane, 2006), we show that the use of ObjectFinder facilitated interactive object search because it combines crucial information about localization (distance and direction) through the detection part with relationships among objects through the description part, as well with the help in navigating to the searched objects.
-However, we have also found that the current system is not efficient enough, is perceived too cumbersome and is not available on the cell phone, and thus cannot be integrated into daily live.
-Our work makes the following key contributions: (1) We make a system contribution and introduce an open-vocabulary interactive object-search prototype that supports location of objects and provision of relevant information.
-(2) We provide an overview of target objects and search challenges in the daily lives of participants under consideration of the potential of assistive technology.
-(3) We critically reflect on the experience and effectiveness of our prototype, and outline opportunities and tensions to be considered in the future development of open-vocabulary interactive object-search systems for real-world application.
+Table 1. Comparison of different systems that can be used for object search.
+System Purpose Enabling Source Interaction Device
+RSA (Eyes, 2024; TapTapSee, Inc., nd; Aira Company, 2024) Multi-Purpose Human Dialogue Smartphone
+ProgramAlly (Herskovitz et al., 2024) Object Search AI Filter Customization Smartphone
+WorldScribe (Chang et al., 2024) Exploration AI Intent Customization Smartphone
+WanderGuide (Kuribayashi et al., 2025) Exploration AI Dialogue, Button-Driven Option Selection Robot
+BeMyAI (Eyes, 2023) Description AI Dialogue Smartphone
+Lookout (Google, 2024) Exploration AI - Smartphone
+Find My Things (Wen et al., 2024) Object Search AI Teachable Object Recognition Smartphone
+LifeInsight (Mathis and Schöning, 2025) Question and Answering AI Dialogue Wearable Device
+ObjectFinder Object Search & Exploration AI Dialogue, Button-Driven Option Selection Wearable Device
+Blind people often face challenges when searching for objects in unfamiliar environments (Müller et al., 2022; Zhang et al., 2021).
+Independently searching for a specific object, such as locating the nearest available chair in a spacious lobby through haptic exploration, can be particularly difficult.
+To search for an object, users would first query the target object, and detect a candidate.
+In order to determine if it is the desired one, both egocentric (e.g.
+11 o’clock, 2.4 meters away”) and allocentric (e.g.next to the desk”) information are necessary for a blind user to perceive the object in their environment (Golledge, 1999; Martolini et al., 2021).
+Upon locating the target object, the user’s intent may vary.
+As shown in examples of Figure 1, once the coffee table is detected, the user might prefer an audio description of items on the table over immediate physical interaction.
+Conversely, if the target is a fan, the user might wish to navigate towards it to turn it on.
+Throughout the search, there may also be a desire to explore the surroundings for better navigation and potentially discover other targets for subsequent searches (Jain et al., 2023b).
+Such an interactive object search task is multifaced, however, no current AI-powered assistive technology can yet handle all the associated subtasks.
+We categorize the existing assistive technologies for blind people into description-based and detection-based systems.
+Description-based systems provide vivid and detailed descriptions of a photo (Eyes, 2023) or brief captions (Corporation, 2024; Lee et al., 2022a; Liu et al., 2023).
+However, these systems are unable to localize a specific object in an unfamiliar environment while ensuring that the target object is in the frame (Xie et al., 2024) (Challenge 1, C1).
+Detection-based systems (Google, 2024; V7, 2024; Zheng et al., 2024), on the other hand, either allow only the search for a limited number of pre-defined objects (Constantinescu et al., 2020a; Yi et al., 2013; Schauerte et al., 2012; Ahmetovic et al., 2020; Kacorri et al., 2017; Google, 2024; V7, 2024; Wen et al., 2024) or provide filtered information (Herskovitz et al., 2024; Chang et al., 2024), limiting understanding of unfamiliar scenes.
+Therefore, when using current detection-based systems in an unfamiliar environment, blind users may not know what is in a room comprehensively and miss items that could be of interest (Challenge 2, C2).
+Apart from that, both detection-based and description-based systems fail to provide egocentric information (distance and direction) or support question and answering directed towards the target.
+The challenges also exist in remote sighted assistance (RSA).
+The procedure by which the remote agents (Aira Company, 2024; Eyes, 2024; TapTapSee, Inc., nd) help to identify objects and describe surroundings involves capturing images from the video feed and zooming in to obtain the necessary visual information (Lee et al., 2020).
+In this context, it is time-consuming for remote agents to adjust the video frame, and they find it challenging to continuously orient the users (Xie et al., 2022a).
+Moreover, recognizing landmarks presents significant difficulties for the agents (Lee et al., 2022b; Xie et al., 2022b; Kamikubo et al., 2020).
+Thus in this work, we aim to address the following question:
+How to integrate the advantages of description- and detection-based assistive systems to support interactive object search by blind people?
+To this end, we designed ObjectFinder, which seamlessly combines open-vocabulary models, an Open-Vocabulary Object Detector (OVOD)(Cheng et al., 2024) and a Multimodal Large Language Model (MLLM)(Achiam et al., 2023), to facilitate an interactive process that ranges from object detection to description for object search.
+Users can input any target via voice commands for object detection, then scan the scene.
+Once a candidate is detected, the system will notify the user to stand still to orient to the target and output real-time egocentric information (distance and direction).
+Following this, users can acquire comprehensive information about their surroundings, tailored to their intent, based on the keyframe captured at the time of detection.
+This process facilitates the identification of potentially interesting and unexpected targets, which can then be explored in further detail during subsequent iterations.
+Based on prior works in object search tasks, we formulated design considerations for an interactive system that enables flexible querying, supports various search subtasks, and adapts the system feedback to user intent.
+We co-designed with a blind person throughout the conception and development of ObjectFinder, across two months and four iterations.
+We deconstruct the complex object search task into the following functions: target object detection, localization, route planning, scene description, and open questions.
+The pipeline for integrating all functions and interaction features was refined based on the iterative feedback from the blind co-designer.
+To evaluate the effectiveness and efficiency of ObjectFinder in facilitating object search, we conducted an exploratory evaluation with eight blind participants.
+They engaged with the system and participated in a semi-structured interview afterwards.
+BeMyAI (Eyes, 2023) and Google Lookout (Google, 2024), popular commercial description- and detection-based systems, were used as baseline comparisons.
+Through thematic analysis (Fereday and Muir-Cochrane, 2006), we demonstrate that the use of ObjectFinder enhanced interactive object search.
+It integrated crucial information about egocentric localization (distance and direction) from the detection component, and allocentric relationships among objects from the description component.
+Additionally, route planning was a valuable feature of ObjectFinder for searching objects.
+Although ObjectFinder provides feedback based on users’ intents, individual variations in procedures and information preferences, also influenced by a scenario’s scope and familiarity, underscore the need for customization and personalization, as discussed later.
+ObjectFinder represents a significant advancement in bridging the technological gap in object search, particularly in unfamiliar contexts.
+Taking advantage of both description- and detection-based systems, its technical approach has the potential to find systems broadly to enhance the independence of blind individuals in their daily lives.
 2. Related Work
-In this section, we provide an overview of existing description-based and detection-based systems for scene understanding and exploration.
-We also discuss the integration of multimodal large language models and interaction features in assistive technologies, which form the background of our study.
-2.1. Description-based and Detection-based Systems for Blind People
-With the advancement of Artificial Intelligence (AI), researchers are developing supplementary AI-based systems to assist blind people in exploring scenes.
-These systems can be classified into description-based and detection-based.
+In this section, we introduce the task of object search for blind individuals and provide an overview of existing description-based and detection-based systems designed specifically for them, which can partially address the task.
+Since no assistive system currently exists for object search in unfamiliar scenarios, we refer to procedures from embodied AI for object search, which typically mimic human behavior.
+This forms the background knowledge for our study.
+2.1. Object Search in Unfamiliar Scenarios
+Object search is a multifaceted task that involves object detection, exploration, navigation, and more.
+In addition to small items that blind individuals frequently search for in their daily lives such as smartphone, keys and wallet (netz-barrierefrei.de, nd), they often search for large, salient objects as landmarks to improve orientation in unfamiliar environments (Project, nd; Yang et al., 2010).
+When searching for objects in unfamiliar environments, blind people typically seek an initial overview of the space, followed by specific details as required (Chang et al., 2024; Shneiderman, 1996).
+If the target object has been found, blind people may have various intentions regarding it.
+For example, they might navigate to the object to interact with it (Herskovitz et al., 2023) (e.g., find a free chair and sit on it), identify a specific object (Brady et al., 2013; Hong and Kacorri, 2024) (e.g., check whether a bottle is shampoo), or perceive the surroundings of the object (e.g., the tabletop (Herskovitz et al., 2023)), which may be too far or inconvenient to touch (Gonzalez Penuela et al., 2024).
+A participant in (Herskovitz et al., 2023) defined the use case of locating an empty chair in the classroom and imagined how the object search system should work: he preferred to scan the environment with smart glasses rather than waving his phone in the crowd, then the system find an empty chair and give directions on how to walk to the chair.
+Some technologies have been proposed to partially address the challenges of object search (Table 1).
+Vizwiz-LocateIt (Bigham et al., 2010) lets users photograph target objects, ask questions to a remote worker on Mechanical Turk, and navigate via sonification.
+Tools such as AIRA (Aira Company, 2024), Vizwiz (Gurari et al., 2018), and BeMyEyes (Eyes, 2024) utilize crowdsourcing to connect blind people with sighted people for real-time remote sighted assistance including object search.
+However, asking the blind people to move their phone to adjust the video frame is time-consuming (Lee et al., 2020).
+WanderGuide (Kuribayashi et al., 2025) has subfunctions for object search implemented on a suitcase, but is designed primarily for exploration without specific consideration of the object search procedure.
+Bhanuka et al.
+(Gamage et al., 2023) suggest that the conversational interface on wearable devices is suitable for the complex task of providing environmental information.
+We categorize existing AI-based assistive technology related to object search into description-based and detection-based systems.
+2.2. Description-based Systems for Blind People
 The description-based systems capture the scene and describe it only once.
-Seeing AI (Corporation, 2024), ImageExplorer (Lee et al., 2022a) and OpenSU (Liu et al., 2023) generate brief image captions for the scenes captured by a mobile phone and enable tactile exploration of the salient objects on the touchscreen.
+Seeing AI (Corporation, 2024), ImageExplorer (Lee et al., 2022a), and OpenSU (Liu et al., 2023) generate brief image captions for the scenes captured by a mobile phone and enable tactile exploration of the salient objects on the touchscreen.
 TapTapSee (TapTapSee, Inc., nd) is an application that generates a concise phrase about the salient object in almost real-time.
-Be My AI, a feature of Be My Eyes (Eyes, 2024) empowered by GPT-4, delivers vivid descriptions of the scenario and allows users to ask questions.
+BeMyAI (Eyes, 2023), a feature of BeMyEyes empowered by GPT-4, delivers vivid descriptions of the scenario and allows users to ask questions.
+Research on BeMyAI (Xie et al., 2024, 2025) indicates that while it serves as a form of distributed cognition, it faces challenges in intent recognition and frequently necessitates the use of multiple images to accurately convey information.
+NaviGPT (Zhang et al., 2025) is a mobile navigation system that provides a brief description of the road ahead.
+LifeInsight (Mathis and Schöning, 2025) is a wearable system embedded with GPT-4 for question answering.
 Some other works focus on the specific features of the salient object, e.g., material (Zheng et al., 2024), transparency (Zhang et al., 2021), and various hazards (Yang et al., 2018a, b).
 However, people with blindness should make the object within the region of interest captured by mobile devices while using description-based applications (Yang et al., 2024; Gurari et al., 2018).
-In addition, we found that people with blindness can barely align the photos they capture with real spatial dimensions using mobile phones during our user study which is in line with the findings of (Kacorri et al., 2017).
+In addition, we found that people with blindness can barely align the photos they capture with real spatial dimensions using mobile phones during our user study, which is in line with the findings of (Kacorri et al., 2017).
+However, Gonzales et al.
+(Gonzalez Penuela et al., 2024) determined that the primary goal for users of AI-powered scene description applications is to identify specific objects.
 Therefore, we implemented an object detector to identify the region of interest for the description module, ensuring a precise understanding of the area where the target is located.
+2.3. Detection-based Systems for Blind People
 Detection-based systems are designed to provide real-time outputs of identified objects or features of interest.
@@
 Various studies have developed wearable systems (Liu, 2023; Islam et al., 2023; Sugashini and Balakrishnan, 2024) with similar functionalities, offering real-time object information through multiple interaction modes.
-Research has explored the detection of personal objects using methods such as SIFT (Constantinescu et al., 2020a; Yi et al., 2013; Schauerte et al., 2012) and advanced deep learning networks (Ahmetovic et al., 2020; Kacorri et al., 2017).
-Navigation assistive systems utilize detection-based methods for obstacle avoidance (Bala et al., 2023; Liu et al., 2021a; Ou et al., 2022), risk assessment (Wang et al., 2024), object finding (Li et al., 2023; Duh et al., 2020; Hu et al., 2022) and passable path planning (Zou et al., 2023; Surougi and McCann, 2023; Hong et al., 2024; Jain et al., 2023).
-These systems autonomously select information, which may limit user agency in actively defining targets.
+WorldScribe (Chang et al., 2024) delivers real-time descriptions of the current view, tailoring the information based on distance and user intent.
+Research has explored the detection of personal objects using methods such as SIFT (Constantinescu et al., 2020a; Yi et al., 2013; Schauerte et al., 2012) and advanced deep learning networks (Ahmetovic et al., 2020; Kacorri et al., 2017; Wen et al., 2024).
+ProgramAlly (Herskovitz et al., 2024) allows users to customize the information filter and efficiently detect specific features of an object.
+This suggests that a predefined list is not preferred for exploration.
+Navigation assistive systems utilize detection-based methods for obstacle avoidance (Bala et al., 2023; Liu et al., 2021; Ou et al., 2022), risk assessment (Wang et al., 2024), object finding (Li et al., 2023; Duh et al., 2020; Hu et al., 2022), shopping (Boldu et al., 2020) and passable path planning (Zou et al., 2023; Surougi and McCann, 2023; Hong et al., 2024; Jain et al., 2023a).
+These systems autonomously select information, which may limit user agency in actively specifying targets.
 Constantinescu et al.
 (Constantinescu et al., 2020b) propose a system that allows users to choose from a limited set of objects to receive audible feedback in their vicinity.
-In detection-based systems, the information available to users is constrained by system limitations, preventing awareness of all existing objects in the environment.
-This study integrates the strengths of description-based and detection-based systems to identify regions of interest and inspire users to unexpected targets.
-Figure 2. Use case diagram : A blind person uses a system to find a socket in an unfamiliar office. The step with dashes is not preferred.
+In detection-based systems, the information available to users is limited by system constraints, which prevents them from gaining a comprehensive understanding of essential scene context for object search.
+2.4. Reference Procedure for Object Search
+Since no existing AI system fully addresses the challenges of object search for blind users, we examine the object-search procedures of embodied AI to guide the design of an assistive object search system.
+Object search is widely recognized as a challenging task that integrates both perceptual and cognitive processes (Sun et al., 2025).
+Typically, embodied agents (Aydemir et al., 2013; Chaplot et al., 2020; Yokoyama et al., 2024) first receive an object query from the user, analyze their surroundings, hypothesize the potential location of the target object, and then plan a navigation path accordingly.
+Recent workflows leveraging LLMs, such as UniGoal (Yin et al., 2025) and SG-Nav (Yin et al., 2024), allow robots to continuously explore their environment and match discovered objects with the intended target.
+CogNav (Cao et al., 2024) investigates the modeling of cognitive process of object search, which involves a broad and contextual search back and forth to build a cognitive map.
+Upon observing the target, it verifies the candidate according to the vicinity, then confirms the candidate.
+Taioli et al.
+(Taioli et al., 2024) equipped the object search agent with a self-questioner and an interaction trigger to produce a refined detection description that includes dialogue regarding the target object.
+In this study, we explore the subtasks involved in object search and integrate them into a unified pipeline specifically designed for blind users, taking advantage of both description- and detection-based systems.
+Figure 2.
+Martin walks into an unfamiliar office and uses an object-search system to search for a socket to charge his smartphone.
+(a) Martin first specifies the target to the system, which then repeats it for confirmation.
+(b) While scanning, candidates are detected.
+The socket “4 meters away at his 10 o’clock next to the trash bin” is not what he wants.
+(c) However, another socket “3 meters away at his 2 o’clock next to the workstation” is the desired one, as he plans to study there.
+(d) After confirming the target, Martin may ask for more details.
+In large rooms, the system should navigate him to the socket.
 […]
-There are eight steps in the diagram: 1) The user tells the system what he/she is looking for (e.g., a socket).
-2)The system gives feedback that it understands what the user wants.
-3)The system indicates it is ready for searching (through a sound).
-4)The user looks around using the system.
-(Function 1) 5a) The system only announces sockets.
-5b) The system continuously informs the user of all objects in the frame (e.g., chairs, table).
-6) Socket detected.
-The system provides feedback with the distance, direction, and relevant points of interest (POls) in the surroundings.
-(Function 2, Function 4).
-7a) The user wants to continue searching.
-The system keeps searching without intervention.
-7b) The user is happy with the found object.
-He/she asks for further details about it, such as directions on how to reach the object and the height of the socket on the wall.
-(Function 5).
-8) For large rooms, the system should navigate the user (e.g., ”Walk around the table on the left side”).
-(Function 3)
+Martin walks into an unfamiliar office and uses an object-search system to search for a socket to charge his smartphone.
+(a) Martin first specifies the target to the system, which then repeats it for confirmation.
+(b) While scanning, candidates are detected.
+The socket “4 meters away at his 10 o’clock next to the trash bin” is not what he wants.
+(c) However, another socket “3 meters away at his 2 o’clock next to the workstation” is the desired one, as he plans to study there.
+(d) After confirming the target, Martin may ask for more details.
+In large rooms, the system should navigate him to the socket.
+3. ObjectFinder
+ObjectFinder is a wearable prototype designed for interactive object search.
+Blind users can specify their target using flexible wording.
+Once the target is detected, they receive real-time egocentric localization information.
+They can further obtain detailed feedback based on their intentions toward the targets.
+We co-designed ObjectFinder with a blind person P0 (see Table 2) by proposing an envisioned scenario, constructing an initial prototype based on it, and then conducting four refinement iterations over two months.
+3.1. Design Goals
+Overall, drawing on related works in object search, we designed ObjectFinder with three primary goals:
+Providing flexibility in target queries and information retrieval.
+Blind people prefer to query flexible target objects and their surroundings, often discovering new items of interest.
+ObjectFinder should facilitate seamless conversational interactions using open-vocabulary models to bridge wording gaps.
+Supporting various subtasks during search.
+Object search is a complex task involving several sequential subtasks: target specification, detection, and feedback generation.
+ObjectFinder should simplify this process by organizing these subtasks into a user-friendly pipeline with accessible interaction features.
+Adapting to the user intent of the target object.
+Blind people exhibit varying intents for identified targets, from navigation to mere perception, and require tailored descriptions or guidance based on the familiarity of their environment.
+ObjectFinder should offer options that allow users to gather system feedback based on their specific intents.
 Figure 3.
 Hardware design and components of the wearable system ObjectFinder.
-It incorporates a stereo camera to capture visual information about the user’s surroundings, a pair of buttons and a microphone to collect the user’s commands.
+It incorporates a stereo camera to capture visual information about the user’s surroundings, a pair of buttons, and a microphone to collect the user’s commands.
 Simultaneously, it executes algorithms through a lightweight processor.
@@
 The smart glasses also include bone conduction earphones for output.
-The image further details the internal components of the glasses which are magnified to show the arrangement from right to left: the right earphone, right infrared camera, color camera, left infrared camera, image ASIC, and left earphone.
+The image further details the internal components of the glasses, which are magnified to show the arrangement from right to left: the right earphone, right infrared camera, color camera, left infrared camera, image ASIC, and left earphone.
 The NVIDIA Jetson Nano, housed in the waist bag, processes the data collected by these components.
-2.2. Multimodal Large Language Models for Assistive Technology for Blind People
-Be My Eyes (Eyes, 2024) is the pioneering application to collaborate with OpenAI, integrating GPT-4 (Achiam et al., 2023) to enable blind individuals to comprehend the real world and interpret images.
-VIALM (Zhao et al., 2024) evaluates the existing Large Language Models (LLMs) for navigation instructions, identifying GPT-4 as the state-of-the-art based on both automatic and human evaluation metrics.
-VIAssist (Yang et al., 2024) identifies six challenges in capturing desirable images for daily requests, such as incomplete target object placement, and offers reshoot suggestions for low-quality images.
-VQAsk (De Marsico et al., 2024) is an Android application that enables interaction with one’s photographs or the surrounding environment through speech and haptic feedback.
-VisionGPT (Wang et al., 2024) leverages LLM to identify existing objects in different scenarios and assess the risk.
-nanoLLaVA (Face, nd) is the existing smallest LLM developed for the deployment on edge devices, e.g.
-assistive technology for blind people.
-Waisberg et al.
-(Waisberg et al., 2024) proposed embedding LLMs in Ray-Ban glasses from Meta to capture the egocentric view of blind people.
-To enable blind people to access visual arts, Zhang et al.
-(Zhang, 2024) designed an interface that enables blind individuals to access visual information from images using LLMs and VLMs.
-GenAssist (Huh et al., 2023) makes image generation accessible by utilizing a sequence of prompts for LLMs.
-In this work, we develop our ObjectFinder by using GPT-4 to generate the description and answer questions.
-In order to generate proper information from LLM, OpenAI offers an extensive tutorial (OpenAI, 2023) on strategies for crafting effective prompts to achieve optimal results.
-Numerous studies have explored various aspects of prompt engineering.
-In this context, we present specific prompt engineering projects designed for the blind community.
-VIALM (Zhao et al., 2024) identifies the importance of prompting LLMs to generate detailed information regarding obstacles and tactile features in navigation instructions.
-Gadiraju et al.
-(Gadiraju et al., 2023) investigate how LLMs can perpetuate subtle yet harmful biases towards communities with special needs, identify specific categories of these biases, and propose training models on diverse resources that positively represent individuals with special needs to mitigate such harms.
-Hao et al.
-(Hao et al., 2024) customize prompts for blind people for various purposes and eliminate any potential offense.
-In this work, we crafted prompts using these strategies to enhance interactive interfaces for blind people, improving digital accessibility and inclusivity.
-2.3. Interaction Features
-Interaction features are crucial for enhancing the usability and effectiveness of assistive technology for blind persons.
-Here we discuss the in- and output modalities separately.
-Speech command input provides a more intuitive and human-like interaction without physical touch, exemplified by voice-controlled smart home devices that offer safety and convenience for blind people (Lee et al., 2022b).
-Voice assistants such as Apple’s Siri and Amazon’s Alexa are widely utilized by the blind community and are deemed beneficial (Oumard et al., 2022b).
-Moreover, advanced navigation systems employ voice commands to define targets and utilize GPS for localization (Hersh, 2022; Ramadhan, 2018).
-He et al.
-(He et al., 2023) enhance the speech command with bone-conducted vibration on head-mounted wearables.
-However, the efficacy of speech understanding systems is contingent upon high accuracy, which can be significantly impacted by environmental noise and the systems’ limited ability to comprehend diverse vocabularies (Oumard et al., 2022a).
-To address these issues, some solutions incorporate simple button-based inputs as alternatives to voice commands (Awad et al., 2018; Tahoun et al., 2020; Hersh, 2022).
-Kamikubo et al.
-(Kamikubo et al., 2024) found that in their study, most users referred to active methods of using buttons, and few users prefer to use voice commands to ask specific questions.
-PathFinder (Kuribayashi et al., 2023) is an automated navigation system with four buttons on the handle interface for functionality selection.
-In our work, we utilize brief voice commands to enable flexible definition of target objects, while employing buttons to streamline the selection of functions.
-In the realm of assistive technologies, methods for conveying information serve distinct purposes.
-Haptic cues, noted for their intuitive reliability, prove effective across indoor and outdoor settings (Bouteraa, 2021).
-Vibration is widely employed for alerts in wearable devices (Mahmud et al., 2014; Babu D et al., 2021; Bouteraa, 2021), while tactile cues (Pielot et al., 2012; Liu et al., 2021b) and kinetic force (Amemiya and Sugiyama, 2008, 2009; Ando et al., 2012) are utilized for indicating navigation directions with handheld devices.
-Yang et al.
-(Yang et al., 2021) explore light cues on hat brims to guide individuals who have light perception.
-Auditory signals, such as beeps of varying frequencies, play crucial roles in warning systems for users and pedestrians (Guerrón et al., 2020; Kayukawa et al., 2019).
-Earcons, which provide contextual information about navigated interfaces, enhance user understanding (Perkins School for the Blind, 2023; OrCam Technologies, 2022).
-Speech remains the predominant output feature on mobile phones (Leporini et al., 2012; Rodrigues et al., 2015) and wearable systems (Bouteraa, 2021; Chen et al., 2021; Boldu et al., 2020; Liu et al., 2021a; Li et al., 2023), offering comprehensive information delivery.
-Notably, bone-conduction devices are preferred for their comfort and wearability, enabling the hearing of system instructions alongside environmental sounds (Asakura, 2021; Liu et al., 2021a).
-Our approach integrates earcons and speech output via bone-conduction headphones within a pair of smart glasses to enhance information accessibility.
-In the remainder of this paper, we outline our three-stage research process for developing our object search prototype: (1) a user-in-the-loop approach for designing and refining the prototype, (2) conducting interviews with additional participants to understand their target objects and challenges related to object search, and (3) performing exploratory user studies in both public and private settings to compare the prototype with existing description-based and detection-based systems.
-3. Phase I: User-in-the-loop Design
-In this phase, we engaged in one-on-one co-design with a blind person (P0, see Table 1) to define functions of the first prototype of the ObjectFinder system, and to develop the design of a study with potential users of the system, rooting initial design decision in the lived experience of a blind person.
-In this phase, we focus on the challenge of combining description-based and detection-based systems as a foundation to address RQ1.
-Here, we provide an overview of our design approach, detail our prototype development process, and present the resulting system.
-Figure 4.
-Left: The initial flowchart based on use case definition.
-The text inside the quotation marks (e.g., “Find”) represents voice input keywords to trigger target identification, target confirmation, and function selection.
-Right: The system architecture includes the interaction steps and multiple functions corresponding to the flowchart.
-Given a task (e.g., “finding a coffee table in an unknown space”), a target with location and orientation information can be found and output.
-The visualization is derived from a real example in our user study.
-[…]
-This figure illustrates the initial flowchart based on the use case definition on the left side and the system architecture on the right side.
-The flowchart has the following elements: 1.
-Vocal input: ’Find ¡target¿’ (block circled 1) a.
-Flows to 2.
-2.
-Confirm ¡target¿?
-a.
-If yes, flows to 3.
-b.
-If no, flows back to 1.
-3.
-Scan the environment, while continuously judging if the target is found.
-(block circled 2) a.
-If yes, flows to 4.
-4.
-Localization information: ¡target¿ ¡distance¿ ¡direction¿ (block circled 3) a.
-Flows to 5.
-5.
-Voice command to choose route planning, scene description, or open questions.
-(block circled 4) a.
-If the voice command ’continue searching’ is given, flows back to 1.
-On the right side of the figure, the network architecture, including the inputs and outputs of YOLO World and MobileSAM, is depicted.
-These components are illustrated to demonstrate their functionality in efficient object detection and enhancing the precision of localization information.
-3.1. Definition of the Use Case
-In a first step, we started with the use case definition, following the principle in (Cockburn, 2000).
-For this purpose, we organized a workshop that included the blind co-designer (P0), one developer, and two accessibility and usability experts (one of them is a blind person).
-The use case search for an object, here a socket, in an unfamiliar environment was identified in the workshop.
-Then, the blind co-designer had the task imagining how he would use a system to look for a socket in an unknown office, while the other participants asked questions that helped the co-designer to define the requirements and the different steps of the search.
-We formulated the use case in a diagram in Figure 2 for a clear illustration.
-The co-designer’s expectation suggested decomposition of the use case into $5$ functions: Object Detection (F1), Localization (F2), Route Planning (F3), Scene Description (F4), and Open Questions (F5).
-Leveraging these categories, we first restructured the interaction features and functions into an initial flowchart, as shown in Figure 4.
-In the following, we explain assembly of hardware components of ObjectFinder, and we outline implementation of basic functionality into the first prototype.
-3.2. Iteration 1: Initial Prototype Design
-In this section, we describe the implementation of the initial prototype based on this use case, detailing the hardware and interaction features and the functionality implementation.
-3.2.1. Hardware and Interaction Features
-Figure 3 presents the system diagram, which comprises a pair of KRVision (KR Vision, [n.
-d.]) smart glasses coupled with a waist bag.
+3.2. Envisioned Scenario
+In an initial step, we began by defining the use case according to the principles outlined in (Cockburn, 2000).
+To achieve this, we conducted a workshop involving the blind co-designer, a developer, and two experts in accessibility and usability, one of whom is blind.
+The co-designer presented a use case for object search: searching for a socket in an unfamiliar office.
+We refined the use case regarding the interaction sequence between the co-designer and the wearable object search system as the envisioned scenario (Carrol, 1995), serves as the basis for our system design, as illustrated in Figure 2 and depicted as follows:
+Martin enters an unfamiliar office, his phone battery depleted.
+In need of power, Martin activates an object search system and commands it to “Find a socket”.
+The system acknowledges the command, and ensures Martin that it has understood the request.
+After Martin confirms, the system signals with a sound, indicating readiness to begin searching.
+As Martin scans the room through the system, he prefers not to be bombarded with information about every detected object; instead, he wants the system to announce only when it detects a socket.
+Upon identifying a socket, the system provides feedback on its egocentric location, including distance and direction, as well as its allocentric relationship with points of interest.
+Using this information, Martin evaluates the suitability of the socket.
+The first socket detected, located near a trash bin about 4 meters away at a 10 o’clock direction, is deemed inconvenient because Martin intends to work at a workstation.
+Therefore, he continues his search for a more suitably placed socket.
+Eventually, a socket near a workstation, just 3 meters away in the 2 o’clock direction, catches Martin’s interest.
+He requests more details about this socket, such as directions to reach it and its height on the wall.
+The system advises Martin to navigate around obstacles, guiding him with instructions like, “Walk around the clutter…”
+Utilizing his cane to detect and avoid clutter, Martin reaches the workstation situated to his front right and successfully charges his phone using the nearby socket.
+3.3. Hardware and Interaction Features
+According to the related work (Herskovitz et al., 2023; Gamage et al., 2023) and the envisioned scenario with the co-designer, a pair of glasses is assumed to be preferred over a smartphone for an object search system.
+We utilize the following hardware to implement this system.
+Figure 3 presents the system diagram, which comprises a pair of KRVision smart glasses (KR Vision, [n.
+d.]) coupled with a waist bag.
 The smart glasses are outfitted with a RealSense R200 RGB-Depth camera, facilitating real-time RGB and depth frame acquisition in an egocentric perspective.
 Additionally, a bone conduction headset is incorporated, enabling auditory output while maintaining perception of environmental sounds.
-In the waist bag, a Jetson Nano, a compact and powerful processor, is utilized for efficient data processing, accompanied by a power bank for energy supplementation.
-The waist bag features two buttons which can be programmed and used for volume adjustment initially.
-A microphone is attached to the collar for audio input.
-As shown in Figure 4, participants simply say commands with keywords to trigger target identification, target confirmation, and function selection.
+In the waist bag, an NVIDIA Jetson Nano, a compact and powerful processor, is utilized for efficient data processing, accompanied by a power bank for energy supplementation.
+The waist bag features two buttons that are programmed for target confirmation and function selection.
+A microphone is attached to the collar for audio input: participants simply speak commands to specify targets or ask questions after triggering open questions.
+The initial version of ObjectFinder is only for the prototype.
+We aim to further integrate all hardware components more compactly to improve the user’s daily experience in real-world use, e.g., with Ray-Ban glasses (Waisberg et al., 2024).
+3.4. Function Implementation
+According to the scenario envisioned in Section 3.2, we have decomposed the object search into five functions: Object Detection (F1), Localization (F2), Route Planning (F3), Scene Description (F4), and Open Questions (F5).
+To integrate these five functions, we define three modules in the pipeline: user specifies the target, system detects the target, and system generates feedback, see Figure 4.
+3.4.1. Module 1: User specifies target.
+When the system is turned on, the frame of the scenario is captured by the smart glasses automatically.
+GPT-4 then generates a list of objects based on the frame to initialize YOLO-World.
+The user specifies the target object using the command frame, “Find ¡target¿”.
+After receiving the command, the system will repeat the target object: “You want to find ¡target¿, please confirm.”.
+For confirmation, the user should press the button with a sticker on the waist bag, while the other button is for respecification.
+Speech-to-text is processed by Google Speech Recognition API, and text-to-speech is handled by the pyttsx3.
+The relationships between the specified target objects and the object classes in YOLO-World are categorized into three types: match, related to, and unrelated to, as shown in Figure 5.
+We define match as instances where the string of one item in the object list appears within the target object.
+This is important because recorded speech may sometimes be unclear due to user rephrasing, such as “Find the chair, no, office chair.” and the surrounding conversations.
+For other cases, the specified target object and each object class in YOLO-World are tokenized and embedded with all-MiniLM-L6-v2 (Reimers and Gurevych, 2021).
+The cosine similarity between the embeddings of the target object and the object classes of YOLO-World is calculated.
+If the similarity score between the target object and any object class in YOLO-World is at least 0.8, a threshold we have set based on empirical data, the target object is deemed to be related to the object class with which it shares the highest similarity in the current list of YOLO-World objects.
+If the target object is deemed related to an existing item in YOLO-World, this item will be used for subsequent object detection searches, thereby eliminating the need to reinitialize the system.
+Conversely, if no similarity between the objects in the list and the target object exceeds the threshold, then the target object is considered unrelated to the current object list of YOLO-World.
+In this case, YOLO-World should be reinitialized with the object list updated to include the new target object.
+During the YOLO-World initialization process, a $3$ Hz beep is played in the background to reassure the user that the system is still operating.
 Figure 4.
-Rectified flowchart.
-Left: The optimized version of the flowchart for the open-vocabulary interactive ObjectFinder.
-Right: Steps added to the ObjectFinder baseline.
-The object detection model in the baseline can only identify a list of objects, the added steps assist in locating the other objects through egocentric scene descriptions.
+ObjectFinder system architecture integrates five functions into three modules for interactive object search.
+(a) Initially, an open-vocabulary object detector, e.g.
+YOLO-World, is initialized with a list of objects extracted from a scenario capture, allowing the user to identify a target object.
+If the target is not on the list, the object detector is reinitialized.
+(b) The user scans the environment.
+If the target is detected, localization information is provided in real-time.
+If not, the user can trigger scene understanding to identify what exists in the scenario.
+(c) The user may activate a sub-branch to obtain further information based on their intent using a multimodal large language model.
+(d) If the user discovers other objects of interest or becomes disoriented, they can reorient themselves to locate the target.
 […]
-On the left side of the figure is the rectified flowchart, and on the right side are the different steps between ObjectFinder and ObjectFinder_Base.
-The rectified flowchart includes the following steps: 1.
-Vocal input: ’Find ¡target¿’ (block circled 1) a.
-Flows to 2.
-2.
-Confirm ¡target¿?
-a.
-If yes, flows to 3.
-b.
-If no, flows back to 1.
-3.
-Scan the environment while continuously judging if the target is found.
-(block circled 2) a.
-If yes, flows to 4.
-b.
-If time out, flows to 5.
-4.
-Localization information: ¡target¿ ¡distance¿ ¡direction¿ (block circled 3) a.
-Flows to 6.
-5.
-Press 1 and 2 to trigger continued searching and scene description.
-(block circled 5) a.
-If 1 is pressed, flows back to 1.
-6.
-Press 1, 2, 3, and 4 to trigger route planning, scene description, continued searching, and open questions.
-(block circled 4) a.
-If 1 is pressed, flows back to 1.
-b.
-Otherwise, conduct the function and flow to 7.
-7.
-Press 1, 2, and 3 to trigger continued searching, repeat, and either scene description or open questions.
-(block circled 6) a.
-If 1 is pressed, flows back to 1.
-b.
-Otherwise, remain in this block.
-On the right, the different steps between blocks circled 1 and 2 are illustrated: 1.
-Confirm ¡target¿?
-a.
-If yes, flows to 2.
-2.
-Judge if the target is in COCO2017.
-a.
-If yes, flows to circled 2.
-b.
-If no, flows to 3.
-3.
-Press 1 to continue searching, press 2 for scene description.
-3.2.2. Functionality Implementation
-To achieve the aforementioned open-vocabulary interaction, all five functions are primarily realized by a combination of models: YOLO-World (Cheng et al., 2024) and MobileSAM (Zhang et al., 2023) as Open-Vocabulary Object Detection (OVOD), GPT-4 (Achiam et al., 2023) as Multimodal Large Language Model (MLLM).
-As the initial flowchart and architecture of the system are shown in Figure 4, we detail the implementation of the individual function in the following.
-Object Detection (F1).
-The recently proposed YOLO-World (Cheng et al., 2024), employing a prompt-then-detect strategy, allows users to define their customized vocabularies for the object detector.
-In our prototype, users specify the target object via a voice command “Find <target>”.
-The prototype then verifies whether the object exists within the predefined vocabulary list.
-If absent, the new target word is embedded and saved in the model’s vocabulary for future reference, a process requiring approximately $20$ seconds for initialization.
-To enhance efficiency, we have pre-populated the vocabulary with a list of common everyday objects.
-We utilize the large version of YOLO-World, which performs inference in $0.5$ seconds on our system.
-The key frame where the object is detected is retained for subsequent analysis.
-Localization (F2).
-Our study utilized an Object-Distance-Direction format for localization as proposed by Constantinescu et al.
-(Constantinescu et al., 2022).
-Directions are presented in a clockwise orientation and distances in meters.
-Initial distances were estimated using YOLO-World’s average bounding box distances due to the pending release of its segmentation version.
-These bounding boxes, which need to overlap ground truth by at least $50\%$, might inaccurately include significant background areas in calculations (Liu et al., 2023).
-To enhance accuracy, MobileSAM (Zhang et al., 2023), the compact version of SAM (Kirillov et al., 2023), was later added to generate segmentation masks using the bounding boxes as prompts post YOLO-World (Cheng et al., 2024).
-Multimodal Large Language Model Functions (F3-F5).
+The image shows a system designed to help blind users search for objects.
+It begins with the user capturing a scene and specifying a target (section a).
+The system initializes object detection using YOLO-World and compares it with the target.
+In the perception module (section b), detected objects are localized, or the system moves to scene understanding if detection times out.
+Based on intent (section c), the system supports navigation (e.g., route planning) or perception tasks (e.g., answering questions or refining the search).
+Feedback loops allow continuous user-system interaction for effective object search.
+3.4.2. Module 2: System detects target.
+Figure 5.
+Initialization with target identification: (a) The list of detectable objects in YOLO-World is initialized with the first capture of the scenario.
+The target objects can be categorized into three types: (b) match, where the object matches an item in the list; (c) related, where the object is related to one item in the list (e.g., “couch” is related to “sofa” with 0.85 similiarity); and (d) unrelated, where the object does not relate to any item in the list.
+In cases where the object is unrelated, the list is updated by adding the target to it.
+[…]
+The figure illustrates how a similarity-based matching process works for object search by blind users.
+On the left, a room scene includes a sofa, cushions, lamp, paintings, and a fan.
+GPT-4 generates an object list from this scene.
+Three example target queries are shown: ”fan,” ”couch,” and ”teapot.” The target ”fan” directly matches the object list, requiring no similarity calculation.
+For ”couch,” the system computes similarities with existing objects: sofa (0.85), cushions (0.61), lamp (0.36), paintings (0.20), and fan (0.27).
+Since ”sofa” has the highest similarity and exceeds the threshold of 0.8, it is selected without needing to reinitialize YOLO-World.
+In contrast, ”teapot” has low similarity scores across all objects—sofa (0.25), cushions (0.25), lamp (0.33), paintings (0.21), and fan (0.19)—none exceeding the threshold.
+As a result, the system triggers YOLO-World reinitialization using the original object list to find the teapot.
+After YOLO-World is initialized with the target object, the user will hear an earcon to signal the start of scanning.
+The system successfully detects the target (F1) when the confidence level of its bounding box exceeds the empirically set threshold of $0.3$.
+The system captures a key frame that includes both RGB and depth information.
+At the same time, another earcon sounds, signaling the user to pause and orient themselves toward the target object.
+This frame is used to calculate localization information (F2) and to query for further intent-based, long-text feedback.
+If the target object is not detected within a time limit of $45$ seconds, it is considered absent in the scenario.
+The user can choose to activate scene understanding (F4) or re-specify the target object.
+If the target is detected successfully, the egocentric localization information will be calculated using the keyframe and delivered in the format Object-Distance-Direction, as proposed by Constantinescu et al.
+(Constantinescu et al., 2022), as illustrated in Figure 6.
+Egocentric information is presented in a clockwise orientation and distances in meters.
+Calculating distance using a bounding box is inaccurate.
+The bounding box for object detection is considered accurate if it overlaps with the ground truth by at least $50\%$.
+However, this criterion might result in the bounding box inaccurately encompassing significant background areas (Liu et al., 2023).
+To enhance accuracy, MobileSAM (Zhang et al., 2023), the compact version of SAM (Kirillov et al., 2023), was later added to generate segmentation masks $M$ using the bounding boxes as prompts post YOLO-World (Cheng et al., 2024).
+The distance of the object is calculated as the average depth from the key frame’s depth map, masked by $M$.
+(1) $Distance=\frac{\sum(\text{depth\_map}\odot M)}{\sum M}$
+To determine the clockwise direction, we use the center of the bottom edge of the frame, denoted as $(x_{c},y_{c})$, as the clock’s center.
+The angle $\theta$ between the center of the bounding box $(x_{\text{bbox}},y_{\text{bbox}})$ and the clock’s center, relative to the bottom edge, is calculated as follows:
+(2) $\theta=\arctan\left(\frac{y_{c}-y_{bbox}}{x_{c}-x_{bbox}}\right)\times\frac{180}{\pi}$
+Then, $\theta$ in the range $(-90^{\circ},90^{\circ})$ is mapped to the clock positions from 9 to 3 o’clock.
+3.4.3. Module 3: System generates feedback.
+When generating feedback, the user is halted by an earcon and oriented towards the target object.
+Simultaneously, the keyframe capturing the user’s egocentric view, which includes the target object, is sent to the MLLM to produce long-text feedback.
+If the user considers the detected candidate to potentially be the target object, they may wish to learn more based on their intent.
+During the refinement iteration with the co-designer, we observed that he primarily had two intents regarding detected objects: navigating to functional objects for interaction, such as finding a charger to charge a smartphone, and perceiving regions of interest, such as identifying items on a tabletop without interacting with them.
+This was followed by a request for further details.
+Therefore, we have currently implemented branches for these intentions in this module, with the possibility of adding more in the future.
+Generating feedback based on user intent.
+In the process of generating intent-based feedback, the system initially uses the keyframe for the first query, while subsequent queries are based on the current egocentric frame of the user.
+In the navigation branch, the user initiates route planning (F3).
+After moving several steps, the target object may no longer be in the field of view, which could prevent further route planning.
+Consequently, the user has the option to either repeat the instruction or trigger a scene description (F4) for orientation.
+For example, if the user is aware that there is a desk on the route to the fan and understands from the scene description (F4) that the desk is directly in front of them, they will know “I’m getting close to the fan.” If the user still feels lost, they can revert to the target specification to relocate the target object.
+In the perception branch, the user can opt to use scene description (F4) to detail the surroundings of the target object, or directly ask open questions (F5) to engage in a conversation about the target object and its surroundings over several rounds.
+As the co-designer always discovered objects of interest or rejected the candidate after learning about the detailed information surrounding it, the user can respecify the target object at any step with ObjectFinder.
+Optimizing interaction features.
+Following the approaches of (Awad et al., 2018; Tahoun et al., 2020; Hersh, 2022), we have programmed the two buttons on the waist bag to select functions.
+We opted for this method over speech commands.
+Despite the flexibility, it is susceptible to environmental noise and the system’s comprehension limitations (Oumard et al., 2022).
+Speech input is used only for target specification and dialogue in open questions (F5) for efficient object search.
+Figure 6.
+Object detection and localization: Each video frame is processed by YOLO-World to detect key frames in which the confidence level of the bounding box around the target object exceeds a certain threshold.
+Subsequently, the segmentation map generated by the bounding box, combined with the depth map, is used to provide precise localization information, including distance and direction.
+[…]
+The figure illustrates a pipeline for object detection and localization.
+A video stream is processed frame by frame using YOLO-World, which identifies objects and highlights keyframes where the target object is detected with high confidence (e.g., a coffee table with a confidence score of 0.88).
+The selected keyframe, including the bounding box, is passed into MobileSAM, a segmentation module.
+Within MobileSAM, the image and bounding box are processed through an image encoder and prompt encoder, respectively, then combined in a mask decoder to produce a segmentation map highlighting the target object.
+This segmented output is then fused with a depth map generated from the corresponding video frame to extract accurate localization information, such as the object’s distance and direction from the camera.
+Enhancing feedback accessibility through prompt engineering.
 According to VIALM (Zhao et al., 2024), GPT-4 is the state-of-the-art for guiding blind people, excelling in both human (correctness, actionability, fluency) and automated evaluations.
@@
 Our system prompt is concrete, incorporating role, tone, and response length, guided by OpenAI’s Prompt Engineering tutorial (OpenAI, 2023).
-Response length, aligned with social media alt-text limits (100-500 characters) (for the Blind, 2024), is set to the maximum to cater to user preferences for vivid responses.
-For navigation instructions, Wang et al.
-(Wang et al., 2023) suggest including position, count, size, color, material, and shape.
-We developed two user prompts: one for route planning (F3) and another for scene description (F4), with voice input for open questions (F5).
-The concrete prompts are provided in the supplementary material.
-3.3. Iteration 2: System Optimization and Refinement
-We conducted four rounds of preliminary half-hour user exploration, inviting P0 to explore the prototype freely in a demo room that integrates two scenarios: a private setting (i.e., a living room) and a public setting (i.e., an office).
-The layouts of these scenarios are shown in Figure 7(b).
-P0 could search for anything and trigger any functions he wanted.
-The wearable prototype is designed to free the hands of blind people to hold a white cane.
-However, it was difficult to determine which information was acquired from our prototype in the small scenarios.
-Therefore, we decided that participants in the user study will not hold a white cane while exploring the system.
-An assistant who is not related to this project should help to ensure the safety of the participants.
-The problems P0 encountered and the improvements of the prototype are summarized in two aspects: interaction features and flowchart.
-The resulting flowchart is formulated in Figure 5.
-3.3.1. Interaction Features
-An acoustic signal recognition module triggers steps like target definition, target confirmation, and option selection in our initial flowchart.
-Due to language redundancy, P0’s voice input wasn’t always fully recognized or understood.
-Inspired by call forwarding, we adapted the system to replace most voice commands with button selections, except for target and open question definitions.
-Consequently, the two buttons on the waist bag, labeled “1” and “2”, handle confirmations and selections: “1” for “yes”, “2” for “no”, and combinations of “1”, “2”, “12”, or “21” for choosing among up to four options.
-We omitted double-clicking due to the buttons’ limited key travel.
-We employed a continuous low-frequency beep to indicate the that system is in operation and enhance user’s sense of safety.
-Additionally, taking cues from Orcam MyEye2 (OrCam Technologies, 2022), we integrated earcons for target detection and input readiness, with beeps at $1Hz$ for system activity and $3Hz$ for model loading and MLLM responses.
-After confirming targets, users should move their heads to scan and stand still following the earcons.
-A misalignment issue between the user’s head and body prompted us to add, “Please align your body with your head’s direction.” to each MLLM output to clarify the expected posture.
+The response length, aligned with the alt text limits of social media (100-500 characters) (for the Blind, 2024), is set to the maximum to accommodate user preferences for vivid responses.
+The co-designer scans the environment by turning his head rather than his body.
+When a target object is captured, he always pauses in his current posture, with his head and body often misaligned.
+Therefore, if the system provides egocentric instructions like “Please walk two steps forward.” or “The desk is in front of you,” it can lead to confusion.
+To resolve this ambiguity, we precede each response from the MLLM with the instruction, “Please align your body with the direction of your head.” This adjustment helps tailor the system’s feedback to accommodate any user’s scanning strategy.
+For route planning, we specifically consider that instructions based on steps are easier for blind people to understand, as suggested by (Gamage et al., 2023).
+Additionally, we provide instructions using landmarks rather than turn-by-turn directions, as per (Jain et al., 2023b).
+The system prompt and two user prompts for route planning (F3) and scene description (F4) are detailed in the supplementary material.
+Table 2. Demographics of participants. P0 is the co-designer who helped to adapt the system to the needs of the target group. P1-P8 were participants of the user study.
 User ID Gender Age Range Vision Level, Onset Experience of Apps
-P0 Male 30-39 Light perception, since about 2004 Be My Eyes, Seeing AI
-P1 Female 20-29 Light perception, since about 2022 Be My Eyes, Seeing AI
-P2 Male 50-59 Fully blind, since birth Be My Eyes, Seeing AI
-P3 Male 20-29 Fully blind, since birth Be My Eyes
+P0 Male 30-39 Light perception, since about 2004 BeMyAI, Seeing AI
+P1 Female 20-29 Light perception, since about 2022 BeMyAI, Seeing AI
+P2 Male 50-59 Fully blind, since birth BeMyAI, Seeing AI
+P3 Male 20-29 Fully blind, since birth BeMyAI
 P4 Male 20-29 Fully blind, since 2010 Lookout
@@
 P8 Male 30-39 Light perception, since 2015 Seeing AI
-Table 1. Participant demographic statistics in Phase II interview with target group.
-3.3.2. Flowchart Refinement
-Participant P0 noted that route planning (F3) was a unique MLLM feature distinguishing our ObjectFinder from Be My Eyes (Eyes, 2024) and frequently activated this feature.
-However, remembering the instructions after a single playback proved challenging, so the instructions needed to be repeated.
-Additionally, P0 might lose orientation after taking several steps toward the target, necessitating additional scene descriptions.
-To address potential confusion regarding the route, we encourage users to search for the target again.
-Consequently, we provided options to continue searching, repeat, and scene descriptions in block 6, following the route planning section in block 4.
-We also introduced open questions in block 6 after scene description (F4) and open question (F5) in block 4 to solicit more detailed feedback.
-In scenarios where the target object was absent, we added block 5 to inform users of available objects.
-(see Figure 5)
-4. Phase II: Interview Study
-During Phase II, we conducted interviews with $8$ participants to understand the technology they use and to identify challenges they face when searching for objects.
-4.1. Method
-To better understand the target objects, we asked participants about the items they most frequently or laboriously search for in their daily lives, and which objects they would be interested in locating using the system, such as furniture or smaller items.
-To gauge whether they require inspiration for unexpected targets (C2), we inquired about the objects they would expect to find in an unfamiliar office or living room, and what items they might not recognize unless informed of their existence.
-We also explored the advantages and disadvantages of the technologies and methods participants currently use for object search or scene understanding.
-Additionally, we specifically asked whether they could match photos taken by a description-based system, such as Be My Eyes, with the actual environment.
-4.1.1. Participants and Procedure
-[…]
-The interview was conducted at the beginning of the exploratory user study described in Sec. 5. The participants were compensated €$40$ for study participation, and travel costs were reimbursed.
-4.1.2. Data Analysis
-We analyzed qualitative data using the thematic analysis framework developed by Fereday and Muir-Cochrane (Fereday and Muir-Cochrane, 2006).
-The first author led the analysis, first reading and re-reading the transcripts, then engaging in coding, and finally organizing the data into themes.
-The themes were subsequently reviewed by the research team.
-In this phase, $122$ data points were categorized into $42$ initial codes, which were then distilled and refined into $15$ codes, from which two main themes were crafted.
-4.2. Results
-Here, we present the two themes crafted from the Phase II interviews, (1) Challenges of Existing Technology for Object Search, and (2) Target Objects and Difficult Cases for Object Search.
-4.2.1. Challenges of Existing Technology for Object Search.
-The data in Table 1 shows that most participants prefer description-based applications.
-Notably, only participant P4 reported using the detection-based application Lookout.
-The use of Be My Eyes for document summarization and volunteer interaction is not included in this table.
-Description-based Systems
-One challenge for blind people is to take the right photo for a description-based system due to the heavy reliance of photography on the visual domain (C1).
-One participant commented on Be My Eyes, “if you lose this point, you can receive the wrong truth.
-A wrong picture you never can be sure, you have to know it if you use these things.
-” (P2).
-Seeing AI and Envision do not help to take the right photo as well, but they do indicate if the photo is of low quality.
-“There is a Seeing AI app for example.
-It tells you, ‘No, this corner is not visible’.
-[…] Envision, it doesn’t tell you unless you already take the photo and it processes the photo, then it tells you ‘no it was quite blurry’.” (P5).
-In terms of learning if the users of Be My Eyes (P0-3) can align the photo with the real environment: all of them answered yes, but participants reported that it is difficult in an unknown environment, e.g., “it’s easier when I know the surroundings, so I can align what’s described.
-But somewhere I’ve never been before, I find it more difficult.”(P1).
-Besides, P2 mentioned a practical method to take good photos: “ you can learn to fix: you can go [to] touch the object, and you go one or two steps back, take your camera and focus.”.
-This suggests that the effectiveness of description-based systems in developing a mental map within an unfamiliar environment without tactile exploration may be limited.
-In addition, participants were suspicious of AI-generated descriptions (Lee et al., 2022a; Xiao et al., 2024) without knowing “when does it start lying?” (P2).
-However, the availability of these applications on the cell phone was considered a significant advantage (P1, P3), while privacy concerns must be considered, as everything is “sent to the cloud” (P0).
-Other methods for Object Search
-As for object-search systems, P2 noted: “at the moment we have no usable object-finding system in most cases.”.
-However, participants mentioned various methods to search for objects.
-For example, P4 uses Lookout, but not frequently due to poor internet connection.
-Some mentioned using tactile exploration and white canes to search for objects in their daily activities (P2, P3, P7).
-As for other technologies, P2 mentioned adding audio descriptions to QR codes with PenFriend to distinguish private items, such as clothing colors, food containers, or medication.
-Besides, participants attached AirTags to their keys, although not everything can be equipped with the technology (P2, P6).
-It enables participants to locate items using sounds.
-For example, they might use an Apple Watch to ping their phone to find their seat in a beer garden.
-However, the aforementioned methods do not have the flexibility to define targets flexibly, and cannot inspire users with interesting unexpected discoveries in unfamiliar scenarios (C2).
-4.2.2. Target Objects and Difficult Cases for Object Search.
-Here, we analyse the role of target objects and difficult cases for object search in both familiar and unfamiliar scenarios.
-Familiar Scenarios
-In their familiar private spaces, the belongings of participants are well-managed and placed in specific locations.
-However, search for them can be difficult if objects are moved by other people (P2, P5), or if participants misplace their own items (P5, P8).
-As P2 mentioned, “ someone was in my office and changed a little thing to put her laptop and something on the table.
-After that, it took an hour to have my things back.”.
-P8 also noted, “ if I’m lazy and put it somewhere, then I really know it has to be around here, but it’s not in its usual place.” Participants also reported searching for small objects in their daily lives: More than one participant had previously looked for their smartphone, keys, small electronics (charger, USB stick, etc.), or items such as coffee cups.
-Sometimes, even though the participants know that objects are nearby, exact location matters, with one participant pointing out that “‘Near’ could be a problem […] your coffee spoiled your laptop.” (P2)
-Unfamiliar Scenarios
-In unfamiliar environments, participants reported difficulties in knowing what is on top of a table (P3, P5, P8) , such as food, because “you usually don’t feel around whole tabletop.” (P3).
-In addition, the participants are unlikely to know the paintings and displays hanging on the wall (P0) and the “hidden” objects (P6).
-Additionally, knowing what one can use in an unknown scenario can be a challenge (C2).
-For example, P4 explained that “we once asked ourselves in the student room whether we had air conditioning.
-[…] The question really occupied us for several hours [and] whether it would be warm in the practice room in the summer.
-[…] If we could scan the room ourselves, we could determine ourselves whether climate regulation is present.” In contrast to the typically small target objects in their daily lives, most participants reported looking for large pieces of furniture when they did not have a mental map of an unknown environment , because “I find it good to get oriented quickly, especially in a foreign room.” (P6).
-P1 typically aimed to find the “next available chair” in an unfamiliar office.
-Meanwhile, P2 sought to locate the cupboard, bed, and other furniture in an unfamiliar hotel room, noting that plugs were harder to identify.
-In smaller rooms, participants commented on the value of tactile exploration to investigate their surroundings, with P7 noting that he “go through the room and feel with my hands to see what I can find next.” However, it is important that this is not feasible in all environments.
-For example, in the specific case of an unfamiliar toilet, “it’s very important to find the objects, because [otherwise] you have to touch everything” (P2).
-Functions Passive Information Aquisition Interactive Object Search
-Lookout (Google, 2024) Be My Eyes (Eyes, 2024) ObjectFinder_Base ObjectFinder
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
 Figure 7.
 Simplified user study procedure focusing on two target objects: a large piece of furniture (an office chair) and a smaller object (a plate of cookies).
-The user walks into an unknown environment, the office in this example.
+The user walks into an unfamiliar environment, the office in this example.
 To sit in front of a desk, he or she should find an office chair first and navigate to it.
@@
 In Step 3, the system provides localization information, indicating ‘office chair found, 2.6 meters, 11 o’clock.’ Step 4 involves route planning with guidance like ‘Six steps forward, follow the edge of the desk, at knee level.’ Step 5 allows the user to sit on the office chair and triggers a scene description: ‘There’s a desk in front of you with a plate of cookies.’ Finally, Step 6 involves prompting the user to ask open questions such as ‘What flavor are the cookies?’, to which the system responds ‘Chocolate!’
-5. Phase III: Exploring User Perspectives on ObjectFinder
-In Phase III, we conducted an exploratory study with blind people to determine if ObjectFinder can support the object search in unfamiliar environments.
-In this step of our work, we include our own prototype ObjectFinder, ObjectFinder_Base (a closed-vocabulary baseline prototype), and the commercial systems Be My Eyes (Eyes, 2024) (a description-based application) and Lookout (Google, 2024) (a detection-based application) as points of reference for participant feedback.
-5.1. Method
-We invited participants P1111-P8888 (Sec.
-5.1.1) to explore the two prototypes and the commercial applications in two unknown environments that were set up in our research lab: a space mimicking a living room, and an office.
-The functional differences between the tested systems are listed in Table 2.
-ObjectFinder_Base differs from ObjectFinder that it only recognizes a limited number of objects with fixed wording.
-The systems are described more detailed in Sec.
-5.1.2.
-In order to assess users’ experiences with the systems and to find out whether an open-vocabulary offers an advantage over a closed-vocabulary approach, we compared ObjectFinder with ObjectFinder_Base in terms of: expectations regarding the functionality, mental map that users formed of the environment after using one system, cognitive load and satisfaction.
-We used semi-structured interviews featuring Likert-scales for function evaluation and open ended questions.
-Questions ranged from “In which areas do you believe our system could be improved?” and “Do you want more or less information?”.
-Additionally, we included the NASA-TLX (Hart, 2006) for assessing the cognitive load to accomplish the tasks, and the Net Promoter Score (NPS) (Reichheld, 2003), a measure that reflects the likelihood if the users would like to introduce the system to their community.
-The exploration of the commercial applications was conducted separately, followed by a short semi-structured interview and the Net Promoter Score (Reichheld, 2003).
-5.1.1. Participants and Procedure
-The participants were the same eight people who took part in the interview described in Phase II (see Sec.
-4.1.1).
-For scene understanding, six of the participants had previous experience using description-based applications such as Seeing AI, Be My Eyes, and Envision, while only one participant uses a detection-based application, Lookout.
-In Table 1, we consider only the scene description or exploration feature of these applications, while usage of the applications for other purposes, such as reading documents, is not included.
+4. User Evaluation
+In order to understand how ObjectFinder can support object search in unfamiliar environments, we conducted an exploratory study with eight blind people.
+In this step of our work, we include our own prototype ObjectFinder, ObjectFinder_Base (a closed-vocabulary baseline prototype), and the commercial systems BeMyAI (Eyes, 2023) (a description-based application) and Lookout (Google, 2024) (a detection-based application) as points of reference for participant feedback.
+Specifically, we focus on the following research questions:
+  1. RQ1:
+    To what extent does ObjectFinder deliver the necessary information for effective object search?
+  2. RQ2:
+    How do blind people perceive the detailed scene context that ObjectFinder provides to facilitate object searches?
+  3. RQ3:
+    How do blind individuals perceive ObjectFinder generally and in comparison to description- and detection-based systems?
+  4. RQ4:
+    What requirements do blind users consider important for an object search system, as their experiences with ObjectFinder show?
+4.1. Participants and Procedure
 […]
-Participants were compensated with €$40$ for study participation, and reimbursed for the travel cost.
-5.1.2. Apparatus
-In this section, we describe the systems we used in the evaluation: ObjectFinder, ObjectFinder_Base, Be My Eyes and Lookout.
-ObjectFinder and ObjectFinder_Base.
-The hardware and flowchart of ObjectFinder and ObjectFinder_Base are identical, except that the baseline only detects object categories in COCO2017 (Lin et al., 2014).
-For ObjectFinder_Base, we set the confidence level threshold to $0.7$ to filter out falsely detected objects.
-In contrast, the YOLO-World used in ObjectFinder can detect any target outside the training dataset, resulting in potentially lower confidence levels.
-We set the confidence threshold for YOLO-World to $0.3$ and tested it with $20$ objects, including those used in our user studies.
-It detected all targets without any false positives, even when the objects were located $8$ meters away.
-Since some target objects are undetectable by the baseline system, we added additional options between blocks 1 and 2, as shown on the right side of Figure 5, to allow participants to trigger scene descriptions multiple times and determine if the target objects are present in the region in front of them.
-Lookout.
-The Explore mode of Lookout, which was tested in this study, detects the nearest object in front of the camera and provides real-time outputs including object categories, quantities, and associated text.
-Be My Eyes. Be My Eyes integrates GPT-4 into its Be My AI feature for a vivid scene description, replacing the need for sighted volunteers.
-Lookout and Be My Eyes reflect the two main parts of our prototype: object detection and scene description.
-However, users cannot actively select their targets, obtain localization information and navigation instructions with these two applications, but instead passively receive information chosen by the system.
-5.1.3. Scenario Exploration
-In each scenario (living room or office), participants were asked to find six target objects: three large pieces of furniture, in order to build a mental map of the unfamiliar environment, followed by three smaller objects found on the desk or coffee table.
+[…]
+4.2. Scenario Exploration
+Blind people typically search for large items as landmarks to construct mental maps of unfamiliar environments, and they would like to use the system to explore small objects on the tabletop.
+In each scenario (living room or office), participants were asked to find six target objects: three large pieces of furniture to establish spatial understanding, followed by three smaller objects found on the coffee table or desk.
 Figure 7 describes a simplified procedure involving the five functions F1-F5.
@@
 Since the closed-vocabulary ObjectFinder_Base can only detect a limited number of target objects (Lin et al., 2014), we categorize the six target objects in each scenario into three groups: unrelated to (two objects), related to (three objects), and exact in (one object) COCO2017.
-As for the related targets, we will hint to the participants to look for the related objects in COCO2017 when they are using the baseline, but they will experience a vocabulary gap between the target definition and MLLM information.
-As for the unrelated objects, the participant cannot even define the target objects (C1).
-So we provide the option that the participant can trigger the scene description function several times to find the objects, as illustrated on the right side of Figure 5.
+As for the related targets, we will hint to the participants to look for the related objects in COCO2017 when they are using the baseline, but they will experience a vocabulary gap between the target specification and MLLM feedback.
+As for the unrelated objects, the participants cannot even specify the target objects (C1).
+So we provide the option for the participants to trigger the scene description function several times to find the objects.
 Cookies are the bonus target, and we observe whether the participants can recognize it themselves to validate the capability of our system to inspire the detection of unexpected targets (C2).
-Figure 8 illustrates several examples of how the participants detected the target objects and received the outputs of the system during the user study.
-The commercial applications, Lookout and Be My Eyes, which were already familiar to the participants and not specifically designed for object search tasks, were briefly used in the living room scenario to locate a fan and a teapot as a way to remind participants of their functions.
+Figure 9 illustrates examples of how the participants detected the target objects and received the system feedback during the user study.
+Lookout and BeMyAI, commercial applications familiar to participants but not designed for object search, were briefly used in the living room scenario to locate a fan and a teapot as reminders of their functions.
 (a) Living Room.
@@
 This figure provides a visual representation of the layout for two scenarios: a living room and an office. It details the sequence and locations of various target objects within these settings.
+Table 3.
+Target objects in each scenario.
+The superscript objectF denotes the function that is activated first upon finding the target objects.
+F3: route planning; F4: scene description; F5: open questions.
 Scenarios Target Objects Names in COCO (Lin et al., 2014) dataset
@@
 Smaller objects teapotF3, bananaF3, flowerF5 ¡None¿, banana, potted plant
-Table 3.
-Target objects in each scenario.
-The superscript objectF denotes the function that is activated first upon finding the target objects.
-F3: route planning; F4: scene description; F5: open questions.
 Figure 9.
-Examples of functions F3-F5 include triggering route planning (F3) to reach both a large piece of furniture and a small item, using scene description (F4) to describe the coffee table, and employing open questions (F5) to inquire about details.
-Postures (first column) and key frames (second column) are visualized, with audio input and output also provided.
+The first column shows examples of postures, while the second column displays key frames captured when target objects were detected (F1) along with the corresponding real-time egocentric localization information (F2).
+The third column presents system feedback generated by the MLLM, which uses route planning (F3) to reach both large and small items, employs scene description (F4) to describe the coffee table, and utilizes open questions (F5) to gather additional details.
 […]
@@
 The pink is more dominant on the edges of the petals, while the yellow is situated toward the center.’
-5.1.4. Data Analysis
-In this phase, we have both qualitative data and quantitative data.
+4.3. Data Analysis
+We have both qualitative and quantitative data.
 For qualitative data, the user study transcripts were analyzed using the hybrid process of inductive and deductive thematic analysis proposed by Fereday and Muir-Cochrane (Fereday and Muir-Cochrane, 2006).
-Analysis was led by the first author who read the transcript repeatedly to achieve familiarisation, followed by multiple rounds of coding.
-Beyond data-driven inductive coding, we also applied deductive coding where it yielded meaningful insights into the system’s capacity to identify regions of interest (C1) and to facilitate the discovery of unexpected targets (C2).
+The first author led the analysis by repeatedly reading the transcript for familiarization and coding it in multiple rounds.
+Beyond data-driven inductive coding, we also applied deductive coding, which yielded meaningful insights into the system’s capacity to identify regions of interest (C1) and to facilitate the discovery of unexpected targets (C2).
 In a workshop, the research team assigned $243$ data points to $69$ codes, which were further refined to $12$ codes, and finally, four themes were crafted and will be presented in Sec.
-5.2.
-Due to the small size of the user group, inferential statistical analysis of the quantitative data is not meaningful. Therefore, we limit our comparison to descriptive statistics.
-5.2. Results
-Here, we present the following four key themes crafted from the interview findings, which focus on the users’ Information Needs for Object Search, their experiences with Mental Mapping, their System Perception and Preferences as well as System Requirements.
-5.2.1. Information Needs for Object Search
-In this section, we analyse the amount of information the participants preferred, and the obligatory and optional information for object search.
+5.
+For quantitative data, we limit our comparison to descriptive statistics due to the small size of the user group.
+5. Findings
+5.1. RQ1: To what extent does ObjectFinder deliver the necessary information for effective object search?
+Participants found that ObjectFinder provides an adequate amount of information, including the obligatory egocentric (distance and direction) and allocentric (relationships among the objects) information, for object search.
+On the other hand, participants have varying perceptions of the optional information (e.g.
+color and alert information).
 Amount of Information
 Regarding the amount of information provided, three participants found it adequate for their needs.
+The average system feedback for route planning (F3) and scene description (F4) contains $62.90$ words, with a standard deviation of $19.11$.
+System feedback for open questions (F5) is relatively shorter.
 However, when asked about their preference for more or less information, responses were divided: half preferred more, while the other half favored less.
@@
 P8 suggested that the information provided could perhaps be reduced after getting the first overview: “it could maybe be a little bit less, so if you know once there are some objects on the table, then you don’t really need this information the second or third time unless you ask the system what’s on the table.” Additionally, we observed that the amount of information varies across functions.
-Be My Eyes, used solely for scene description, received praise from three participants for its “detailed” information, although one deemed it excessive.
+BeMyAI, used solely for scene description, received praise from three participants for its “detailed” information, although one deemed it excessive.
 In contrast, participants noted that our ObjectFinder delivered more information than necessary for efficient navigation.
@@
 I think sometimes meters might be more useful […] So it sometimes has like $0.1$ meter […] you could have said ‘right in front of you’ or ‘one step’ like that.” Participants mentioned that the distances reported were inaccurate, appearing larger than they perceived.
-Upon reviewing the video, we observed that some discrepancies were caused by measurements taken from the head to the object, rather than horizontally, as shown in Figure 11.
+Upon reviewing the video, we observed that some discrepancies were caused by measurements taken from the head to the object, rather than horizontally.
 This inaccuracy may be more pronounced in our small indoor settings.
@@
 Contextual information around the target is crucial for object search, such as relationships among surrounding objects and their location information.
-For example, P3 noted, “it’s good that the objects around it are announced, so you have some idea where the object you want to find is in relation to other objects.” A description should also include the distance which says something about the user as a reference point to the object.
-P8 suggested to improve ObjectFinder by incorporating distance into the MLLM output, “ when you make the description, it holds everything but it does not really talk about the distance.”
+For example, P3 noted, “it’s good that the objects around it are announced, so you have some idea where the object you want to find is in relation to other objects.” A description should also include the distance, which says something about the user as a reference point to the object.
+P8 suggested to improve ObjectFinder by incorporating distance into the MLLM output, “ when you make the description, it holds everything, but it does not really talk about the distance.”
 Optional Information
@@
 Scores for the first three functions exceeded 4, while the remaining two stayed below 4.
-Figure 10. Net Promoter Scores (NPS) for different systems, assessed by participants P1-P8 while exploring scenarios.
-[…]
-The bar chart of the net promoter scores for different systems: ObjectFinder (M=6.75, SD=2.66), ObjectFinder_Base (M= 5.88, SD=2.42), ObjectFinder with efficient hardware (M=8.13, SD=2.01), Be My Eyes (M=7.25, SD=3.01), Lookout (M=5.31, SD=2.43).
-Figure 11.
-Special cases.
-Upper half: Initially, the participant did not detect the coffee table as an obstacle to the sofa.
-Subsequently, the participant stepped backward to detect the coffee table instead of lowering the head.
-Bottom Left: The participant needed cues to lower the head to detect items on the desk.
-Bottom Right: The participant felt uneasy about the distance ($1.5$ meters) to the office chair because the measurement was taken from the head to the chair, not horizontally.
-[…]
-This visualization depicts the postures and key frames for special cases where participants needed cues to lower their heads, and felt uncertain about the distance.
-In the example shown in the lower right part of the figure, a participant who is about 1.7 meters tall detects a chair in her vicinity (0.5 meters away), yet the distance is calculated to be 1.5 meters from the glasses to the chair.
-5.2.2. Mental Mapping
+5.2. RQ2: How do blind people perceive the detailed scene context that ObjectFinder provides to facilitate object search?
+Participants were able to gain an overview of the scene, orient themselves, and explore search options within the detailed scene context provided by ObjectFinder.
 There is some evidence that participants obtained more detailed environmental information through the use of ObjectFinder.
 P1 noted: “I think it was very good [to know] where I am.
-When it said where my target object was, it also told me the surrounding objects, because maybe if I find the surrounding object first, then I know, OK I’m close.
+When it said where my target object was, it also told me the surrounding objects, because maybe if I find the surrounding object first, then I know, OK, I’m close.
 ” P6 noted the benefits of discovering other usable items (C2) and understanding their arrangement, “I didn’t know there was another chair there.
-It gives more information about objects and their arrangement.” Participants wished to get an overview of unfamiliar rooms without receiving too many details.
-P8 explained, “I could imagine if you really don’t know the room, and you want to first [have an] overview what is in the room, it was pretty detailed.
+It gives more information about objects and their arrangement.” Participants would like to get an overview of unfamiliar rooms without receiving too many details.
+P8 explained, “I could imagine if you really don’t know the room, and you want to first [have an] overview of what is in the room, it was pretty detailed.
 ”
@@
 “I found it easier to think about which direction.” (P6).
-This highlights the potential of ObjectFinder to support users in building a mental map of an unknown space.
-This was more difficult using the existing commercial systems.
-For example, after using Be My AI, P8 mentioned, ”Be My AI had no information regarding the distances, so you have […] a less sense of location ”.
-As previously analyzed, localization information (distance and direction) and the relationships between surrounding objects are crucial for effective object search.
-The lack of localization information (P2, P8) and the absence of context regarding object relationships (P2) were mentioned as reasons why participants did not prefer Lookout, highlighting the potential for solutions that can combine multiple aspects of object search.
-5.2.3. System Perception and Preferences
-In this section, we analyze the general perception of ObjectFinder, the user preference with the systems, and the cognitive load to finish the tasks.
+5.3. RQ3: How do blind individuals perceive ObjectFinder generally and in comparison to description- and detection-based systems?
+Participants generally prefer ObjectFinder for object search over description- and detection-based systems because it enables users to specify targets actively and provides both egocentric and allocentric essential information, though some practice is required to become accustomed to it.
 General Perception of ObjectFinder
@@
 ObjectFinder was also appreciated for accurate object searches without physical touch.
-Subsequently, participants rated the system’s helpfulness ($\mu=4.13,\sigma=0.641$) and their independence ($\mu=4.31,\sigma=1.01$) with ObjectFinder in an unfamiliar environment, with both ratings based on $5$-point Likert scales and averaged over Through ObjectFinder, participants discovered more than just target items, finding inspiration in unexpected objects, both small and large (C2).
+Subsequently, participants rated the system’s helpfulness ($\mu=4.13,\sigma=0.641$) and their independence ($\mu=4.31,\sigma=1.01$) with ObjectFinder in an unfamiliar environment, with both ratings based on $5$-point Likert scales and averaged over $4$.
+Through ObjectFinder, participants discovered more than just target items, finding inspiration in unexpected objects, both small and large (C2).
 For example, “I’m sort of inspired because now I know what options I have, instead of walking in and just looking for my phone because I know it must be there, but I don’t know what else could be there […] I was inspired to look for, for example, bananas and stuff like that.”.
@@
 User Preferences and Cognitive Load
-The NPS for the prototypes does not vary substantially between the scenarios, office ($\mu=6.25,\sigma=2.49$) and living room ($\mu=6.38,\sigma=2.67$), with scores being generally high, and not lagging behind those allocated to commercially available systems (see Figure 10).
-With respect to differences between the systems and implications for user preferences, participants appreciated that they could actively define the target objects with ObjectFinder, making it more reliable in searching for specific items.
+Regarding the differences between the commercial applications and the prototypes, participants valued that they could actively define the target objects with ObjectFinder, making it more reliable in searching for specific items.
 As P6 mentioned, “what I also liked about this system is speaking to it, I find that more targeted.”.
 P3 echoed this sentiment, highlighting the biggest advantage: “The biggest advantage […] contrary to BeMyAI is that you can specify what you want to find.” (C1).
+”Furthermore, as previously mentioned, ObjectFinder delivers essential information for object searches that cannot be provided by either description-based or detection-based systems.
+This was more difficult using the existing commercial systems.
+For example, after using BeMyAI, P8 mentioned, ”BeMyAI had no information regarding the distances, so you have […] a less sense of location ”.
+As previously analyzed, egocentric localization information (distance and direction) and the allocentric relationships among surrounding objects are crucial for effective object search.
+The lack of localization information (P2, P8) and the absence of context regarding object relationships (P2) were mentioned as reasons why participants did not prefer Lookout, highlighting the potential for solutions that can combine multiple aspects of object search.
 As shown in Table 4, the study also examined cognitive load, showing that it was relatively low across system types and scenarios.
-Retrieving keyframes for MLLM information and participant feedback reveals that the landscape orientation of the camera on the glasses, contrasting with participants’ usual practice of taking portrait photos using commercial applications, and the presence of relatively low furniture in the living room often result in the prototype failing to recognize the coffee table as an obstacle to the sofa, as illustrated in Figure 11.
+Retrieving keyframes for MLLM information and participant feedback reveals that the landscape orientation of the camera on the glasses, contrasting with participants’ usual practice of taking portrait photos using commercial applications, and the presence of relatively low furniture in the living room often result in the prototype failing to recognize the coffee table as an obstacle to the sofa.
 The exploratory experience with the prototype was not comprehensive enough.
 P7 mentioned that F4 and F5 are insufficiently tested in our study, and added: “one needs to get used to it, it requires some practice, and then the search will be very precise.” P5 mentioned that she was not accustomed to the earcon indicating to stand still.
+Table 4.
+Assessment of workload required to complete tasks with two systems (ObjectFinder vs.
+ObjectFinder_Base) and in two scenarios (Office vs.
+Living Room) using NASA-TLX.
+The scores range from 1 (very low) to 21 (very high).
 Sub-scale Systems Scenarios
@@
 Frustration 4.13 3.48 4.38 3.07 4.00 3.46 4.50 3.07
-Table 4.
-Assessment of workload required to complete tasks with two systems (ObjectFinder vs.
-ObjectFinder_Base) and in two scenarios (Office vs.
-Living Room) using NASA-TLX.
-The scores range from 1 (very low) to 21 (very high).
-5.2.4. System Requirements
-In this section, we analyze the system requirements, i.e. interaction features, software and hardware requirements.
+5.4. RQ4: What requirements do blind users consider important for an object search system, as their experiences with ObjectFinder show?
+Participants basically suggested that efficiency is crucial in terms of interaction features and hardware for object search.
+Additionally, they considered object detection, localization, and route planning to be important functions of object search.
 Interaction Features.
@@
 P1 preferred pressing the button without waiting for the entire list to be read aloud.
-Additionally, P8 described the voice commands as “pretty, pretty easy”, while P4 expressed concerns with programs that solely accept voice input, suggesting “maybe it’s better to have the ability to switch between a list of maybe recognized objects and voice commands.” Besides the current earcons, P6 suggested creating an additional earcon for option confirmation, rather than using spoken text.
+Regarding target specification, P8 described the voice commands as “pretty, pretty easy”, while P4 expressed concerns with programs that solely accept voice input, suggesting “maybe it’s better to have the ability to switch between a list of maybe recognized objects and voice commands.” Besides the current earcons, P6 suggested creating an additional earcon for option confirmation, rather than using spoken text.
 As mentioned before, some participants preferred more information, while they suggested “implementing a skip option” (P1) and the ability to “switch information on and off optionally” (P2).
@@
 We asked participants to rate the functions before and after using ObjectFinder and ObjectFinder_Base in the explored scenarios, in terms of their perceived importance and level of interest, as captured by Likert scores (Figure 10).
-From the ratings, the functions object detection, localization information and route planning are important for the object search task ($\mu{>}4.0$), accompanied by a reduced standard error.
+From the ratings, the functions object detection, localization information, and route planning are important for the object search task ($\mu{>}4.0$), accompanied by a reduced standard error.
 Among the five functions evaluated, localization information was rated as the most important.
@@
 We note that the software should be fast and make fewer mistakes.
-P2 mentioned that receiving explanations from a sighted person on Be My Eyes “works better because it’s without delay.” Half of the participants mentioned frequent misidentifications of Lookout was not preferred.
+P2 mentioned that receiving explanations from a sighted person on BeMyEyes “works better because it’s without delay.” Half of the participants mentioned that frequent misidentifications of Lookout were not preferred.
 As P5 noted, “it misidentified objects […], like there’s no dishwasher,” and at times, the teapot was categorized as a helmet or mouse.
 Hardware Requirements
-Five participants mentioned that the hardware was heavy for long-term use.
-We asked participants to imagine that the ObjectFinder prototype was implemented on more efficient hardware (Waisberg et al., 2024), such as the Ray-Ban glasses for image capturing paired with an iPhone for data processing, while actually trying on a pair of Ray-Ban glasses to evaluate their feel.
-With these enhancements, the Net Promoter Score (NPS) of ObjectFinder increased from ${\mu}{=}6.750,{\sigma}{=}2.659$ to ${\mu}{=}8.125,{\sigma}{=}2.013$, as illustrated in Figure 10.
-In comparison, a common advantage of the applications, Be My AI and Lookout, is the availability on the smartphone.
 Participants noted the distinct advantages and disadvantages of capturing scenarios with glasses compared to a cellphone.
@@
 6. Discussion
-In this section, we summarize our key findings to address our research questions, discuss potential use cases for the object-search system, and reflect on requirements to be addressed by future work for object-search systems to be of use to people who are blind.
-RQ1: How can an open-vocabulary prototype integrate detection-based and description-based technologies to support object search for blind people, and what challenges are associated with its design and development?
-Overall, our experience with the design and implementation of ObjectFinder suggests that it is feasible to create assistive technology for object search that combines detection-based and description-based technologies, and that can be leveraged by blind people wishing to examine unfamiliar environments.
-In particular, our work shows that such a system can facilitate exploration of different settings (i.e., living and office environments), and that it can be a first step toward building a mental map of an unknown space.
-In particular, the description part of the ObjectFinder prototype offered an overview that helped participants understand the environment, know what to expect in the scenario, and discover unknown objects for subsequent searches.
-Compared to the description-only and detection-only applications, systems with the design of ObjectFinder have potential to provide a guiding function and the comprehensive information necessary for object search: localization information (distance and direction) and the relationship among the surrounding objects.
-However, our findings also show that there are several design challenges that remain unaddressed.
-For example, open questions regarding hardware design and interaction features remain, and addressing these is crucial when creating technology that people are interested in integrating into their daily lives.
-With respect to the information provided by the system, results suggest that the amount of information and whether to make additional details optional for object search, such as color and alert information, needs to be fine-tuned based on participants’ information processing strategies.
-RQ2: What are users’ experiences and preferences regarding the open-vocabulary interactive object-search system, specifically comparing the description-based and the detection-based, with the hybrid ObjectFinder prototype?
-Given the absence of existing technology to assist blind individuals in object search, most participants expressed enthusiasm for this open-vocabulary interactive object-search prototype.
-Participant ratings reflect satisfaction with the prototypical integration of functions within ObjectFinder, and suggest that cognitive load during system use was acceptable.
-In comparison with exclusively description-based or detection-based systems, our prototype effectively defines the target object, locates regions of interest, and delivers essential localization information and route planning, with interview results showing that this was appreciated by users.
-With respect to user experience with specific features of ObjectFinder, our work shows that detection of small objects was well-supported, whetheras the user experience when exploring larger space was less smooth, with open questions regarding user safety remaining (i.e., better obstacle recognition).
-6.1. Future Directions for the Development of Object-Search Systems
 Our research highlights that object-search systems have potential as an assistive technology for people who are blind.
-Here, we want to highlight future directions for the development of such systems, outlining challenges and opportunity for the Human-Computer Interaction and accessibility research communities.
-6.1.1. Potential Features for Future Integration
+Here, we want to discuss future directions for the development of such systems, outlining challenges and opportunity for the Human-Computer Interaction and accessibility research communities.
+6.1. Potential Features for Future Integration
 In our study, we uncovered a range of features and characteristics of ObjectFinder that offer interesting avenues for future work, either through iteration on our system, or integration in other, comparable systems.
@@
 Providing descriptive overviews of unfamiliar environments is essential for revealing unexpected objects and should be relative to the user’s position.
-Based on our findings, we recommend to include both distance and direction in descriptions, with far objects quantified in meters and nearer objects described in steps or as ‘right in front of you’.
+Based on our findings, we recommend to include both distance and direction in descriptions, with far objects quantified in meters and nearer objects described in steps or as “right in front of you”.
 These details can facilitate room exploration, help in building mental maps, and assist in orientation.
 Understanding advantages and drawbacks of additional information.
-Although the essential information for object search, localization information and relationships among surrounding objects, is well-defined, participants had mixed feelings about additional details like color and alert information.
+Although the essential information for object search, localization information, and relationships among surrounding objects, is well-defined, participants had mixed feelings about additional details like color and alert information.
 Consequently, the future object-search system should allow users to select the amount of information by incorporating options to skip information or to stop and continue information output.
@@
 On the other hand, this may address concerns with respect to social accessibility, i.e., the visibility of assistive technology to others, and associated stigma (Shinohara and Wobbrock, 2016).
-6.1.2. Tensions and Concerns Regarding Vision- and AI-based Assistive Technology
-There are tensions and concerns that need to be resolved for systems such as ObjectFinder effectively and safely supporting object finding.
+6.2. Tensions and Concerns Regarding Vision- and AI-based Assistive Technology
+There are tensions and concerns that need to be resolved for systems such as ObjectFinder to effectively and safely support object searching.
 User habits and needs regarding lighting conditions for camera-based systems.
-With respect to technical requirements of current approaches to computer vision systems, we note that lighting conditions are crucial for camera-based systems.
+With respect to the technical requirements of current approaches to computer vision systems, we note that lighting conditions are crucial for camera-based systems.
 However, this conflicts with the fact that light plays a different role in the lives of blind people: Legally blind persons may not use light in their homes and workplaces in the same way as (typically) sighted system developers would anticipate  (Gonzalez Penuela et al., 2024), and people who do have residual vision may not find lighting conditions required by camera systems comfortable.
@@
 With advancements in wearable systems and smart glasses, researching and integrating these specific scanning techniques into mobility training is both viable and beneficial.
-Moreover, incorporating sensors and cameras with wider angles, like LiDAR (Liu et al., 2021a) and omnidirectional camera (Kawaharazuka et al., 2024), could enhance scene perception over larger areas, reducing the need for physical scanning efforts.
+Moreover, incorporating sensors and cameras with wider angles, like LiDAR (Liu et al., 2021) and omnidirectional camera (Kawaharazuka et al., 2024), could enhance scene perception over larger areas, reducing the need for physical scanning efforts.
 However, despite potential technological solutions (which may come with new challenges), this specific instance highlights tensions around safety: On the one hand, users are invited to rely on systems for object detection, on the other hand, it is known that vision- and AI-based systems can be unreliable in specific situations (e.g., in the context of autonomous driving (Cummings and Bauchwitz, 2024)), and even more so in the context of disability (e.g., see (Kane et al., 2020)).
@@
 Understanding the limitations of technology for object search.
-Finally, the Human-Computer Interaction and accessibility communities have previously discussed issues surrounding technologies focusing on independence of users.
+Finally, the Human-Computer Interaction and accessibility communities have previously discussed issues surrounding technologies focusing on the independence of users.
 In particular, Vincenzi (Vincenzi, 2021) contributed a critical appraisal of assistive technology for navigation of blind people, suggesting that there were instances where working with other people was more relevant than applying a technology-based solution.
@@
 Here, we need to ask critical questions around specific features of ObjectFinder, and we want to leave you with one example: Is it really necessary for a blind person to use the system in the workplace, or could non-disabled colleagues make a bigger effort to not misplace or alter their desk?
-7. Limitations and Future Work
-There are a few limitations that need to be taken into account when interpreting our findings.
-This research, conducted in small indoor settings of our lab, does not investigate the effectiveness of ObjectFinder in larger, real-world public, private, or outdoor environments.
-We therefore reached only participants from nearby regions.
-The requirement for a WiFi connection by ObjectFinder, due to its use of GPT-4, limits its operational range and may introduce security concerns akin to those observed with Be My Eyes.
-Although ObjectFinder possesses MLLM capabilities to infer surroundings beyond the immediate view, it lacks the ability to precisely remember explored spaces, rendering every space unfamiliar to its current version and thus providing known information to the user.
-Additionally, while participants preferred our ObjectFinder when paired with efficient hardware (Waisberg et al., 2024), we only had them try the hardware and imagine its implementation, as it was not actually deployed on more efficient hardware.
-We see several avenues for future work.
-With the development of smart glasses, we will pursue collaborations with companies to create a more comfortable and affordable system, which will integrate ObjectFinder’s functionality to support object search and build up mental maps.
-After addressing the challenges and implementing the suggestions from participants, we plan to conduct a larger-scale quantitative evaluation of the system with blind individuals in real private and public scenarios across different cities.
-This approach would provide an opportunity to engage a larger number of participants and to evaluate the effectiveness of ObjectFinder in real-world scenarios, which we initially opted not to do at scale.
-Our first goal was to assess whether the open-vocabulary interactive object-search system addresses existing challenges in daily life and whether our implementation delivers an independent and confident experience to the target group in controlled settings.
 7. Conclusion
@@
 With our prototype, we address shortcomings of existing systems that are either description- or detection-based: locating regions of interest and discovering incidental targets.
-Results of the exploratory user study suggest that the approach is promising, offering blind people a tool to develop a better understanding of unfamiliar spaces.
-Overall, our work represents an initial step toward developing AI-based assistive technology that supports object search, providing first insights into user requirements and application challenges.
+The system feedback is tailored to various user intents, and our exploratory user study suggests that this approach is promising, as it provides essential egocentric localization and allocentric scene context while enabling interactive object search.
+Overall, our work represents an initial step towards developing AI-based assistive technology that supports object search, providing first insights into user requirements and application challenges.
 Here, we hope that our work will encourage and facilitate further development of object-search systems, and that it will inspire future studies into the experiences that blind people have with such technologies.